### v1

# mathx”00 mathx”08 mathx”10 mathx”18 mathx”20 mathx”28 mathx”30 mathx”38 mathx”40 mathx”44

Abstract
We present a construction of the finite-volume massive sine-Gordon model in the UV subcritical regime using a renormalization group method. The resulting measure has Gaussian tails, respects toroidal symmetries and is reflection-positive.
1 Introduction
1.1 Initial remarks
The sine-Gordon model is quite popular in the statistical mechanics and quantum field theory literature. For an early study see [Col75]. The first rigorous renormalization group study of the model began with the Roman school in the series of papers [BGN82, Nic83, NRS86, RS86]. Soon thereafter [BK87] studied the model by a direct analysis of the Polchinski equation, but could not cover the full subcritical regime. With [BY90] the foundation of an alternative method was laid. [DH93] was the first attempt at controlling the full subcritical regime in this approach, but unfortunately an error spoiled the results. The later [DH00] also contains an error. [^1] This was only spotted recently by the author. The issue is that (56) does not follow from Lemma 6 as claimed. It is probably possible to fix the proof by switching back to the more local norms used in [DH93], but we are not yet in a position to adjudicate whether this would cause any other problems. The present work is an alternative solution which avoids the problematic cluster expansion step altogether. This does come at a price: without a cluster expansion, it is not clear how to implement the usual leveraging of the mass for good control of the infinite volume behavior. The next substantial development was the injection into the RG framework of the finite-range decomposition [MS00, BMS03]. Since then there has been further investigation of this and closely related models, especially in the critical phase (cf. e.g. [Fal12, Fal17, BPR24, BPR24a]).
The present work is primarily a technical note: to the best of our knowledge, in the available literature there is currently no complete construction of the model with Brydges-Yau-type methods in the full UV-subcritical (or superrenormalizable) regime, even in finite volume (this might be considered a folklore result). We also take this opportunity to establish some additional properties of the limiting measures. It should be possible, perhaps with substantial modifications of the method, to extend (at least some of) our results to infinite volume, which question we hope to revisit in the future.
Our main references were [DH00, SS09, Fal12]. If anything is ever unclear, it is a good bet that it is covered in one of these works.
1.2 Structural remarks
This paper is designed to be read as follows. The main numbered sections constitute a minority of the text, since there are many technical notions and results that we have relegated to the appendices. When we think it is time for the reader to familiarize themselves with the contents of a given appendix, we make a note of this in the body of the text. From that point onward, it is assumed that the reader understands the given appendix and can independently refer to it when they feel the need to refresh their memory.
1.3 The model
We study the sine-Gordon measures. Let $\beta\in[4\pi,8\pi)$ and $\zeta$ be a bounded Borel function on the unit torus $\mathbb{T}^{2}$. [^2] The restriction on $\beta$ away from 0 is not necessary, we have made it to avoid constants depending on $\beta^{-1}$ which seems to only be a distraction. The interested reader can easily modify the results herein to cover this case, which can anyway be treated by other methods giving stronger results. Furthermore pick $\varepsilon>0$ and then let $\gamma[C_{\varepsilon}]$ be the law of the mollified unit mass GFF $\chi_{\varepsilon}*\Phi$ (where $\chi_{\varepsilon}$ is a standard mollifier). [^3] We put in a mass to keep things as simple as possible, primarily to avoid issues with the zero mode which arise in the massless model. Setting it equal to $1$ is convenient because it equals the width of the torus. Of course, this is not a “generic” choice, but the additional complications in case the mass is not fixed to the torus size did not seem illuminating enough to warrant investigation in this note. Put
[…]
$E[\varepsilon,\zeta]$ is chosen to make $\mu_{\mathrm{sG}}$ a probability measure and the double brackets denote Wick ordering:
[…]
$\mu_{\mathrm{sG}}$ does make sense because $\mathrm{C}_{\varepsilon}$ is a UV-mollified GFF covariance with cutoff $\varepsilon$ so that the Gaussian measure $\gamma[\mathrm{C}_{\varepsilon}]$ is $\mathcal{C}^{\infty}$-valued. If $\mathrm{T}$ is a symmetry of the torus,
$\mathrm{T}_{*}\mu_{\mathrm{sG}}[\varepsilon,\zeta]=\mu_{\mathrm{sG}}[\varepsilon,\mathrm{T}\zeta].$ (3)
At this point we fix some constants. First, $\sigma\coloneqq 1-\frac{\beta}{8\pi}\in(0,\frac{1}{2}]$ is the distance to criticality. $\delta$ is a small positive number which could be fixed at, say, $\delta=\frac{1}{100}$ throughout, $\ell\gtrsim_{\sigma^{-1},\delta^{-1}}1$ is an integer, $L=\ell^{\mathfrak{m}}$ where $\mathfrak{m}=\ell^{2}$, $A\gtrsim_{L}1$, and $\zeta_{*}\lesssim_{A^{-1}}1$. We also introduce $\mathfrak{c}^{(1)},\mathfrak{c}^{(2)},\mathfrak{c}^{(3)},\mathfrak{c}^{(4)}$ with $\mathfrak{c}^{(2)}\lesssim_{1}1\lesssim_{1}\mathfrak{c}^{(1)}\wedge\mathfrak{c}^{(3)}$, $\mathfrak{c}^{(4)}\gtrsim_{1}\mathfrak{c}^{(3)}$, and $\mathfrak{h}_{*}\gtrsim_{1}1$, $\kappa_{*}=\mathfrak{c}^{(4)}\log L$. $\zeta_{\bullet}$ is a completely arbitrary positive number and will represent the upper bound on the size of the coupling $\zeta$ as measured in $\mathrm{L}^{\infty}$. $j_{\bullet}$ is a negative integer defined as
$j_{\bullet}=\bigg{\lfloor}\sigma^{-1}\log_{L}\!\bigg{(}\frac{\zeta_{*}}{\zeta_{\bullet}}\bigg{)}\bigg{\rfloor}\wedge 0$ (4)
and will serve as the terminus for the UV RG iteration; it is chosen so that if $\|\zeta\|_{\mathrm{L}^{\infty}}\leqslant\zeta_{\bullet}$, the effective coupling at scale $j_{\bullet}$ is bounded by $\zeta_{*}$. Write $\mathrm{B}_{\bullet}=\zeta_{\bullet}\,\overline{\mathrm{Ball}}(\mathrm{L}^{\infty})$ topologized as a subset of $\mathrm{L}^{\wp}$ where $\wp=5\sigma^{-1}$ (anything strictly larger than $4\sigma^{-1}$ would work); it is a closed subset. Finally, we introduce a maximal mollification scale
$\varepsilon_{\bullet}=\frac{1}{6}L^{j_{\bullet}-1}.$ (5)
An important convention we will follow is that $\mathcal{O}_{x_{1},\ldots,x_{n}}$ denotes some constant depending on the specified parameters and this can change from line to line, which spares us from naming the various transient quantities that appear in estimates and cannot be comfortably absorbed with the $\lesssim$ notation.
The Laplace transforms for $J\in\mathcal{C}^{\infty}(\mathbb{T}^{2})$ are
$\mathrm{Z}[\varepsilon,\zeta](J)=\int e^{\langle J,\phi\rangle}\mu_{\mathrm{sG}}[\varepsilon,\phi](\mathrm{d}\phi).$ (6)
These are Fréchet-analytic functionals on $\mathcal{C}^{\infty}(\mathbb{T}^{2})$. To measure their size, we introduce the space $\mathfrak{C}$ of continuous functionals on $\mathcal{C}^{\infty}(\mathbb{T}^{2})$ with finite norm
$\|F\|_{\mathfrak{C}}=\sup_{J\in\mathcal{C}^{\infty}(\mathbb{T}^{2})}e^{-\gamma\|J\|^{2}_{*}}|F(J)|;$ (7)
here $\gamma$ is $\mathcal{O}_{1}$ and $\|\!\cdot\!\|_{*}$ is a Sobolev norm with sufficiently large regularity (e.g. $10$ derivatives will certainly suffice). We also define the subset $\mathfrak{L}\subset\mathfrak{C}$ of those functionals which are Laplace transforms of a probability measure. It turns out that $\mathfrak{L}$ is closed, convergence in $\mathfrak{L}$ implies weak convergence of the corresponding probability measures on $\mathcal{D}^{\prime}(\mathbb{T}^{2})$, and these measures all have Gaussian tails (see Appendix E). Now we have everything necessary to state the central result.
Proposition 1.1.
The map $\mathrm{Z}[\cdot,\cdot]:(0,\varepsilon_{\bullet}]\times\mathrm{B}_{\bullet}\longrightarrow\mathfrak{C}$ is Lipschitz.
The point of this proposition is that $\mathrm{Z}[\cdot,\cdot]$ automatically extend to the completion $0,\varepsilon_{\bullet}]\times\mathrm{B}_{\bullet}$, i.e. we get a UV limit. We also write $\mathrm{Z}[0,\zeta]$ for the corresponding limiting objects. Note that $\zeta_{\bullet}$ was arbitrary, so we in fact get this result for all couplings, but the Lipschitz constant is not uniform, so we have found this the most natural formulation of the result.
The analysis we do should also carry through to small complex coupling. This would show that $\mathrm{Z}[0,\zeta]$ have analytic extensions in $\zeta$ around $\zeta=0$, which implies that the usual perturbation theory is convergent, a non-trivial result (though we emphasize that it is much less clear if in the complex case, we get corresponding complex-valued measures; this is because Levy’s theorem for complex measures is more delicate). We have elected to restrict just to real coupling to avoid more complications regarding the finer points of infinite-dimensional holomorphy.
The “toroidal” OS axioms follow:
Corollary 1.2.
Let $\zeta$ be constant. The probability measure $\mu_{\mathrm{sG}}[0,\zeta]$ is reflection-positive and invariant under toroidal symmetries. If $0<\zeta\leqslant\zeta_{*}$, it is also non-Gaussian.
Proof.
The covariance property (3) transfers by weak convergence to the limiting measures and so
$\mathrm{T}_{*}\mu_{\mathrm{sG}}[0,\zeta]=\mu_{\mathrm{sG}}[0,\mathrm{T}\zeta]=\mu_{\mathrm{sG}}[0,\zeta].$ (8)
For reflection positivity, view $\mathbb{T}$ as $(-1/2,1/2]^{2}$ and define $A_{n}$ to be the subset $(-1/2,1/2]\times[n^{-1},1/2-n^{-1}]$. Let $\zeta_{n}=\zeta\,\mathbb{I}_{A_{n}\cup\vartheta A_{n}}$ be the coupling which equals $\zeta$ on $A_{n}\cup\vartheta A_{n}$ ($\vartheta$ being the reflection) and vanishes otherwise. As $\zeta_{n}\rightarrow\zeta$ in $\mathrm{L}^{\wp}$, it suffices to establish RP for the approximating measures $\mu_{\mathrm{sG}}[0,\zeta_{n}]$. Let $F(\phi)=\sum_{i}c_{i}e^{\langle J_{i},\phi\rangle}$ be an exponential functional which depends only on the field values in $(-1/2,1/2]\times[\varepsilon^{\prime},1-\varepsilon^{\prime}]$ (i.e. $\mathrm{supp}(J_{i})\subset\mathbb{R}\times(-1/2,1/2]\times[\varepsilon^{\prime},1-\varepsilon^{\prime}]$). Then if $\varepsilon<\varepsilon^{\prime}$,
[…]
$\geqslant 0$ (9)
where the bound follows by the reflection positivity of $\mathrm{C}$. By taking $\varepsilon\rightarrow 0$, since
$\langle\vartheta F,F\rangle_{\mathrm{L}^{2}(\mu_{\mathrm{sG}}[\varepsilon,\zeta_{n}])}=\sum_{i,j}c_{i}\overline{c_{j}}\,\mathrm{Z}[\varepsilon,\zeta_{n}](J_{i}+\vartheta J_{j})$ (10)
[…]
Non-Gaussianity follows from the detailed structure of the Laplace transform, see Section 2.4. ∎
2 Renormalization group flow
2.1 Progressive integration
We explicate a method to iteratively compute the integral that defines $\mathrm{Z}[\varepsilon,\zeta]$, giving us the ability to analyze these Laplace transforms sufficiently accurately to obtain the stated results. In this section, we just sketch the basic ideas, deferring to later sections for precise justification.
$\mathrm{C}_{\varepsilon}$ admits a scale decomposition into $\Gamma_{\hskip-1.0pt\varepsilon;j}$ with corresponding fields $\Psi_{\hskip-1.0pt\varepsilon;j}$. Refer to Appendix A. Therefore by shifting the integration,
$\mathrm{Z}$ $[\varepsilon,\zeta](J)$
[…]
Now we define
$\zeta_{\varepsilon;j}=L^{2j}e^{\frac{1}{2}\beta\sum_{k=j}^{0}\Gamma_{\hskip-1.0pt\varepsilon;k}(0)}\mathrm{S}_{L^{-j}}\zeta,\quad\text{i.e.}\quad\zeta_{\varepsilon;j+1}=L^{2}e^{-\frac{1}{2}\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\mathrm{S}_{L^{-1}}\zeta_{\varepsilon;j}.$ (12)
The flow equation for $j\leqslant-1$ is
$\mathcal{Z}_{-\mathsf{N}-1}[\varepsilon,\zeta](\phi)$ $=\exp\!\bigg{(}\int_{\mathbb{T}_{-\mathsf{N}-1}^{2}}\zeta_{\varepsilon;-\mathsf{N}-1}\cos\!\big{(}\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\phi\big{)}\bigg{)},$
$\mathcal{Z}_{j+1}[\varepsilon,\zeta](\phi)$ $=\mathbb{E}\big{[}\mathcal{Z}_{j}(\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}.$ (13)
By induction, it is easily verified that
$\mathrm{Z}[\varepsilon,\zeta](J)=e^{\frac{1}{2}\langle\mathrm{C}_{\varepsilon}J,J\rangle+E[\varepsilon,\zeta]}\mathbb{E}\bigg{[}\mathcal{Z}_{j}[\varepsilon,\zeta]\bigg{(}\sum_{k=j}^{0}\mathrm{S}_{L^{k-j}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j}}\mathrm{C}_{\varepsilon}J\bigg{)}\bigg{]}.$ (14)
2.2 The ansatz
To actually solve the iteration, we need a good ansatz, which is provided by polymer activities. Refer to Appendices B.1, B.3, B.4. At this point we do the algebraic manipulations which however are not yet justified since we do not know if the objects involved have enough integrability, so we simply assume that they do. This will be verified once we introduce appropriate norms. Anyway, with $K_{j}$ a polymer activity,
$\mathcal{Z}_{j}[\varepsilon,\zeta](\phi)$ $=e^{\mathcal{E}_{j}[\varepsilon,\zeta]}\sum_{X\in\mathcal{P}_{j}}e^{V_{j}[\varepsilon,\zeta](X^{c},\phi)}\prod_{Y\in\mathcal{C}(X)}K_{j}[\varepsilon,\zeta](Y,\phi),$
$V_{j}(X,\phi)$ $=\int_{\mathbb{T}_{j}^{2}}\zeta_{\varepsilon;j}\cos\!\big{(}\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\phi\big{)}.$ (15)
Now we compute the RG map. That is, we want to compute $\mathbb{E}\big{[}\mathcal{Z}_{j}[\varepsilon,\zeta](\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$. We wish to represent $\mathcal{Z}_{j+1}$ as
$\mathcal{Z}_{j+1}[\varepsilon,\zeta]=e^{\mathcal{E}_{j+1}[\varepsilon,\zeta]}\sum_{X\in\mathcal{P}_{j+1}}e^{V_{j+1}[\varepsilon,\zeta](X^{c},\phi)}\prod_{Y\in\mathcal{C}(X)}K_{j+1}[\varepsilon,\zeta](Y,\phi).$ (16)
First put
$\tilde{K}_{j}[\varepsilon,\zeta](X,\phi,\psi)=\mathbb{I}_{X\in\mathcal{P}_{c}}\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X}e^{V_{j}[\varepsilon,\zeta](LX-Y,\mathrm{S}_{L}\phi+\psi)}\prod_{Z\in\mathcal{C}(Y)}K_{j+1}[\varepsilon,\zeta](Z,\mathrm{S}_{L}\phi+\psi).$ (17)
An important point is that there is some ambiguity in the representation of $\mathcal{Z}_{j}$: terms can be moved out of $K_{j}$ and into $\mathcal{E}_{j}$. We will need to make use of this ambiguity to track some “bad” terms by removing them from $K_{j}$ before evaluating the expectation, otherwise $K_{j}$ would grow uncontrollably along the flow. This step is usually called extraction. To perform it, we take an activity $Q_{j}[\varepsilon,\zeta](X)$ which is nonzero only on small $X$ and does not depend on the field. [^4] Field dependence is the generic situation in other models; it poses no further issues by itself. The resulting fact that the counterterms in this model are field-independent is nice because it does not spoil the duality with a Yukawa or Coulomb gas. We also write $\delta\mathcal{E}_{j}[\varepsilon,\zeta](X)=\sum_{B\in\mathcal{B}_{j}(X)}\delta\mathcal{E}_{j}[\varepsilon,\zeta](B)$ for an activity which will correct $\mathcal{E}_{j}$ in a way which compensates the subtraction of $Q_{j}$. Finally, as an intermediate quantity, define for any $Y\in\mathcal{S}_{j+1}$ and $B\in\mathcal{B}_{j+1}(Y)$
$J_{j}[\varepsilon,\zeta](Y,B)=\sum_{D\in\mathcal{B}_{j}(LB),X\in\,\mathcal{S}_{j}}^{D\subset X,\bar{X}^{L}=Y}\frac{1}{|X|}Q_{j}[\varepsilon,\zeta](X)-\mathbb{I}_{Y=B}\sum_{Y^{\prime}\in\mathcal{S}}\sum_{D\in\mathcal{B}_{j}(LB),X\in\,\mathcal{S}_{j}}^{D\subset X,\bar{X}^{L}=Y^{\prime}}\frac{1}{|X|}Q_{j}[\varepsilon,\zeta](X).$ (18)
Note that this vanishes unless $B\subset Y\in\mathcal{S}$. The second term is only there to ensure
$\sum_{Y\in\mathcal{S}}J_{j}[\varepsilon,\zeta](Y,B)=0.$ (19)
Another identity that $J_{j}$ satisfies is
$\sum_{B\in\mathcal{B}_{j+1}(Y)}J_{j}[\varepsilon,\zeta](Y,B)=\sum_{X\in\mathcal{S}_{j}}^{\bar{X}^{L}=Y}Q_{j}[\varepsilon,\zeta](X)-\mathbb{I}_{|Y|=1}\sum_{D\in\mathcal{B}_{j}(LY),X\in\mathcal{S}_{j}}^{D\subset X}\frac{1}{|X|}Q_{j}[\varepsilon,\zeta](X).$ (20)
For notational convenience, if $(Z,\mathfrak{B})$ is a pair of a polymer and a choice $\mathfrak{B}=(\mathfrak{B}_{Y})_{Y\in\mathcal{C}(Z)}$ of a square $\mathfrak{B}_{Y}\in\mathcal{B}_{j+1}(Y)$ in every connected component of $Z$, we write
$J_{j}^{\mathfrak{B}\rightarrow Z}[\varepsilon,\zeta]=\prod_{Y\in\mathcal{C}(Z)}J_{j}[\varepsilon,\zeta](Y,\mathfrak{B}_{Y}).$ (21)
Finally, for $X\in\mathcal{P}_{j+1}$ define
$P_{j}^{X}[\varepsilon,\zeta](\phi,\psi)$ $=\prod_{B\in\mathcal{B}_{j+1}(X)}\big{(}e^{V_{j}[\varepsilon,\zeta](LB,\mathrm{S}_{L}\phi+\psi)}-e^{V_{j+1}[\varepsilon,\zeta](B,\phi)+\delta\mathcal{E}_{j}[\varepsilon,\zeta](LB)}\big{)},$ (22)
$R_{j}^{X}[\varepsilon,\zeta](\phi,\psi)$ $=\prod_{Y\in\mathcal{C}(X)}\bigg{(}\tilde{K}_{j}[\varepsilon,\zeta](Y,\phi,\psi)-\sum_{B\in\mathcal{B}_{j+1}(Y)}J_{j}[\varepsilon,\zeta](Y,B)\bigg{)}.$ (23)
From now we omit the arguments $\varepsilon,\zeta$ if context permits. Plugging all these definitions into the formula for $\mathcal{Z}_{j+1}$ gives
$\mathcal{Z}_{j+1}(\phi)=e^{\mathcal{E}_{j}}\sum_{X_{1},X_{2},Y_{1},\mathfrak{B}\rightarrow Y_{2}}^{*}e^{V_{j+1}(X_{1},\phi)+\delta\mathcal{E}_{j}(LX_{1})}\mathbb{E}\big{[}P_{j}^{X_{2}}(\phi,\Psi_{\hskip-1.0ptj})R_{j}^{Y_{1}}(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}.$ (24)
The *-sum is over quadruples in $\mathcal{P}_{j+1}^{\times 3}\times(\mathfrak{B}\rightarrow\mathcal{P}_{j+1})$ for which $X_{1}\cup X_{2}\cup Y_{1}\cup Y_{2}=\mathbb{T}_{j+1}^{2}$, all of these sets have disjoint interiors and $Y_{1},Y_{2}$ are completely disjoint. We make one last adjustment to exhibit linear cancellations. Define $X=X_{2}\cup Y_{1}\cup\bigcup\mathfrak{B}^{*}$. Then $X_{1}=X^{c}\cup(X-X_{2}\cup Y_{1}\cup Y_{2})$ and we find the final form
$\mathcal{Z}_{j+1}(\phi)=e$ $^{\mathcal{E}_{j}+\delta\mathcal{E}_{j}(\mathbb{T}_{j}^{2})}\sum_{X\in\mathcal{P}_{j+1}}e^{V_{j+1}(X^{c},\phi)}\cdot$ (25)
$\prod_{X^{\prime}\in\mathcal{C}(X)}\sum_{(Z,Y_{1},\mathfrak{B}\rightarrow Y_{2})\rightarrow X^{\prime}}e^{V_{j+1}(X^{\prime}-Z\cup Y_{1}\cup Y_{2},\phi)-\delta\mathcal{E}_{j}(L(Z\cup Y_{1}\cup Y_{2}))}\mathbb{E}\big{[}P_{j}^{Z}(\phi,\Psi_{\hskip-1.0ptj})R_{j}^{Y_{1}}(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}.$
The expectations have partially factorized due to the finite-range property. The sum in the second line is over triples in $\mathcal{P}_{j+1}^{\times 2}\times(\mathfrak{B}\rightarrow\mathcal{P}_{j+1})$ with $Z,Y_{1},Y_{2}$ disjoint and $X=Z\cup Y_{1}\cup\bigcup\mathfrak{B}^{*}$. We remark that $Y_{2}=X$ is impossible since it would force $X\subsetneq\mathfrak{B}_{W}^{*}$ for at least one $W\in\mathcal{C}(Y_{2})$. Anyway, we now pin down the definition of the RG flow.
Definition 2.1.
Let $(\varepsilon,\zeta)\in(0,\varepsilon_{0}]\times\mathrm{B}_{\mathrm{cs}}(\zeta_{*})$. The RG flow is the dynamical system trajectory
$(\zeta_{j},K_{j},\mathcal{E}_{j})_{j=-\mathsf{N}-1}^{0}$ (26)
with evolution ($X$ connected)
$\zeta_{j+1}$ $=L^{2}e^{-\frac{1}{2}\beta\Gamma_{\hskip-1.0ptj}(0)}\mathrm{S}_{L^{-1}}\zeta_{j},$ (27)
$K_{j+1}(X,\phi)$ $=\sum_{(Z,Y_{1},\mathfrak{B}\rightarrow Y_{2})\rightarrow X}e^{V_{j+1}(X-Z\cup Y_{1}\cup Y_{2},\phi)-\delta\mathcal{E}_{j}(L(Z\cup Y_{1}\cup Y_{2}))}\mathbb{E}\big{[}P_{j}^{Z}(\phi,\Psi_{\hskip-1.0ptj})R_{j}^{Y_{1}}(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}},$ (28)
$\mathcal{E}_{j+1}$ $=\mathcal{E}_{j}+\delta\mathcal{E}_{j}(\mathbb{T}_{j}^{2}).$ (29)
We remark that, because $Q_{j}$ does not depend on $\phi$, $K_{j}$ depends only on $\phi$ within $X$. Also, the initial step simplifies drastically:
$K_{-\mathsf{N}}(X,\phi)$ $=\mathbb{E}\big{[}P_{-\mathsf{N}-1}^{X}(\phi,\Psi_{\hskip-1.0pt-\mathsf{N}-1})\big{]},$ (30)
$\mathcal{E}_{-\mathsf{N}}$ $=0.$ (31)
Anyway, to better understand what this map is doing, we look at the linearization around the trivial RG flow when $\zeta=0$ (this is the flow of the GFF). We assume that $\delta\mathcal{E}_{j}$ and $Q_{j}$ will be linear in $K_{j}$. Then the linearization is
$\mathcal{L}K_{j}(X,\phi)=\sum_{Y\in\mathcal{P}_{\mathrm{c};j}}^{\bar{Y}^{L}=X}\big{(}\mathbb{E}\big{[}K_{j}(Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj}\big{]}-Q_{j}(Y)\big{)}+\mathbb{I}_{|X|=1}\sum_{B\in\mathcal{B}_{j}(LX)}\bigg{(}\sum_{Y\in\mathcal{S}_{j}}^{B\subset Y}\frac{1}{|Y|}Q_{j}(Y)-\delta\mathcal{E}_{j}(B)\bigg{)}.$ (32)
To obtain this, we have used (19) and also the identity $\mathbb{E}\big{[}V_{j}(LX,\mathrm{S}_{L}+\Psi_{\hskip-1.0ptj})\big{]}=V_{j+1}(X,\phi)$, which is the reason for our choice of $\zeta_{j}$ in the first place. We can define $\delta\mathcal{E}_{j}$ to remove the last term and then we may choose $Q_{j}$ to cancel some terms in $K_{j}$.
To identify these problematic terms, we first do a Fourier analysis in the zero mode of the field. That is, note that $K_{j}(X,\phi+2\pi\beta^{-1/2})=K_{j}(X,\phi)$ (this is clearly preserved by the flow); then put
$k_{j}^{(q)}(X,\phi)=\frac{1}{2\pi}\int_{-\pi}^{\pi}e^{-iqx}K_{j}(X,\phi+\beta^{-1/2}x)\,\mathrm{d}x.$ (33)
We will see that the kinds of norms preserved by the flow in fact imply that $K_{j}(X)$ is analytic on a strip (refer to Appendix B.4.1), so $k_{j}^{(q)}$ are well-decaying in $|q|$ and we also have
$K_{j}(X,\phi)$ $=\sum_{q\in\mathbb{Z}}k_{j}^{(q)}(X,\phi),$ (34)
$k_{j}^{(q)}(X,\phi+x)$ $=e^{i\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}qx}k_{j}^{(q)}(X,\phi)$ (35)
with the second making sense even for complex $x$ with $\mathrm{Im}(x)$ not too large. In any case, the extraction $Q_{j}$ is taken from the neutral ($q=0$) terms:
$Q_{j}(X)=\mathbb{I}_{X\in\mathcal{S}}\mathbb{E}\big{[}k_{j}^{(0)}(X,\Psi_{\hskip-1.0ptj})\big{]}.$ (36)
Having thus retained control of the RG up to the last step, we are left with handling $\Gamma_{\hskip-1.0pt0}$ which is slightly different but ultimately easier.
This concludes the basic exposition of the setup. In the next section we start working up to analytic estimates on the RG flow.
2.3 Analytic control
To study the RG flow, we wish to control
$\|K_{j}[\varepsilon,\zeta]-K_{j}[\dot{\varepsilon},\dot{\zeta}]\|_{\mathfrak{h},\kappa,A}.$ (37)
That is, we are comparing the fluctuation step where we allow the fluctuation covariance to change; this is important because we want to compare different flows to obtain convergence.
We will always assume $\dot{\varepsilon}\leqslant\varepsilon$ with corresponding $\dot{\mathsf{N}}\geqslant\mathsf{N}$, and we will mostly deal with $j\geqslant\mathsf{-N}$. Since the case $j=-\mathsf{N}-1$ has slightly worse estimates, we need to treat it separately in principle, but we have tried to make the presentation as seamless in this regard as we could manage; there are only some qualifications in the propositions below to deal with this border case at the specific parameter values with which it actually arises.
Since there will be a lot of estimates, we make some notational abbreviations. We write
$V_{j}=V_{j}[\varepsilon,\zeta],$ $\quad\dot{V}_{j}=V_{j}[\dot{\varepsilon},\dot{\zeta}],$
$K_{j}=K_{j}[\varepsilon,\zeta],$ $\quad\dot{K}_{j}=K_{j}[\dot{\varepsilon},\dot{\zeta}],$ (38)
and similarly for $J_{j}[\varepsilon,\zeta]$, $P^{X}_{j}[\varepsilon,\zeta]$, etc. In case a mixed argument appears, such as $K_{j}[\varepsilon,\dot{\zeta}]$, we will not use an abbreviation. Regarding norms, if $x_{1},\ldots,x_{n}$ belong to the same normed space, write
$\|x_{1},\ldots,x_{n}\|=\max_{i=1,\ldots,n}\|x_{i}\|.$ (39)
Regarding the various parameters, we assume $\mathfrak{h}\in[\mathfrak{h}_{*},2\mathfrak{h}_{*}]$ and $\kappa\in[\kappa_{*},2\kappa_{*}]$.
As a general rule of thumb, an estimate on a quantity such as $\|K_{j}-\dot{K}_{j}\|_{\ldots}$ should reduce to an estimate of the form $\|K_{j}\|_{\ldots}$ simply by taking $\dot{\zeta}=0,\dot{\varepsilon}=\varepsilon$, which kills $\dot{K}_{j}$ and any possible dependence on $\varepsilon$. This is not quite correct because in the general difference estimates, the right-hand side will sometimes involve $2\mathfrak{h}$ whereas the left involves only $\mathfrak{h}$, while this is not the case in the non-difference estimate, so speaking precisely, one has to derive the latter separately; but we will not write this out, since the derivation is always a simpler version of the difference estimate. As a sanity check, the reader can consider the next proposition and reason whether the difference estimates do reduce to their non-difference versions in this way.
Proposition 2.2.
Take $\mathsf{N}\leqslant j<j_{\bullet}$. The potential satisfies
$\|V_{j}-\dot{V}_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\lesssim_{1}|X|\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{)},$ (40)
$\|e^{V_{j}}-e^{\dot{V}_{j}}\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\lesssim_{1}2^{|X|}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{)}.$ (41)
Assuming furthermore that we have inductively established $K_{j},\dot{K}_{j}\in\mathbb{B}^{j,1}_{2\mathfrak{h},2\kappa,A}$,
$|Q_{j}(X)-\dot{Q}_{j}(X)|$ $\lesssim_{1}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle},$ (42)
$|\delta\mathcal{E}_{j}(B)-\delta\dot{\mathcal{E}}_{j}(B)|$ $\lesssim_{1}\max_{B\subset X\in\mathcal{S}}\big{(}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}\big{)},$ (43)
$|J_{j}^{\mathfrak{B}\rightarrow X}-\dot{J}_{j}^{\mathfrak{B}\rightarrow X}|$ $\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-2|X^{*}|}\|K_{j},\dot{K}_{j}\|^{|\mathcal{C}(X)|-1}_{\mathfrak{h},\kappa,A}\cdot$
$\hskip 20.0pt\big{(}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)},$ (44)
$\|P_{j}^{(\cdot)}(\cdot,\psi)\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\leqslant\mathcal{O}_{L}^{|X|}\big{(}\|\zeta_{0;j}\|_{\wp}+\|K_{j}\|_{\mathfrak{h},\kappa,A}\big{)}^{\!|X|},$ (45)
$\|P^{(\cdot)}_{j}(\cdot,\psi)-\dot{P}^{(\cdot)}_{j}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\leqslant\mathcal{O}_{L}^{|X|}\big{(}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}^{\!|X|-1}\cdot$
$\hskip 20.0pt\Big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle(LX)^{*}\rangle}+$
$\hskip 30.0ptL^{-j}|\varepsilon-\dot{\varepsilon}|\big{(}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}+$
$\hskip 40.0pt\|\psi-\dot{\psi}\|_{\mathcal{C}^{2}(LX)}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\Big{)},$ (46)
$\|R_{j}^{(\cdot)}(\cdot,\psi)\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-(1+\eta/3)|X|}G_{\kappa}(LX,\mathrm{S}_{L}\phi+\psi)\|K_{j}\|_{\mathfrak{h},\kappa,A}^{|\mathcal{C}(X)|},$ (47)
$\|R_{j}^{(\cdot)}(\cdot,\psi)-\dot{R}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-(1+\eta/3)|X|}\sup_{t\in[0,1]}\!G_{\kappa}(LX,\mathrm{S}_{L}\phi+(1-t)\psi+t\dot{\psi})\cdot$
$\hskip-50.0pt\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}^{|\mathcal{C}(X)|-1}\Big{(}\big{(}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}+$
$\hskip 70.0pt\|K_{j},\dot{K}_{j}\|_{2\mathfrak{h},\kappa,A}\big{(}L^{-j}|\varepsilon-\dot{\varepsilon}|+\|\psi-\dot{\psi}\|_{\mathcal{C}^{2}(LX)}\big{)}\Big{)}.$ (48)
The bounds for $P^{(\cdot)}_{j}$ also hold in the case $\varepsilon=\dot{\varepsilon}$, $\psi=\dot{\psi}$, and $j=-\mathsf{N}-1$.
Note that we have merely assumed $K_{j},\dot{K}_{j}\in\mathbb{B}^{j,1}_{2\mathfrak{h},2\kappa,A}$, though we will in fact be able to get $K_{j},\dot{K}_{j}\in\mathbb{B}^{j,2-\delta}_{2\mathfrak{h},2\kappa,A}$. Since the proofs involved are somewhat lengthy, we move them to Appendix D.
Now we assemble all these estimates into a bound on $\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}$. We also move this proof into Appendix D.
Proposition 2.3.
Take $\mathsf{N}\leqslant j<j_{\bullet}$. Suppose $K_{j}\in\mathbb{B}^{j,1}_{\mathfrak{h},\kappa,A}$. Then
$\|K_{j+1}\|_{\mathfrak{h},\kappa,A}\leqslant\mathcal{O}_{1}L^{2\sigma}\|K_{j}\|_{\mathfrak{h},\kappa,A}+\mathcal{O}_{A}\|\zeta_{0;j}\|^{2}_{\wp}.$ (49)
Now suppose $K_{j}\in\mathbb{B}^{j,1}_{2\mathfrak{h},2\kappa,A}$. Then
$\|K_{j+1}$ $-\dot{K}_{j+1}\|_{\mathfrak{h},\kappa,A}\leqslant$
$\mathcal{O}_{1}L^{2\sigma}\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}],K_{j}[\dot{\varepsilon},\zeta]-K_{j}[\dot{\varepsilon},\dot{\zeta}]\|_{\mathfrak{h},\kappa,A}+$
$\mathcal{O}_{A}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}],K_{j}[\dot{\varepsilon},\zeta]-K_{j}[\dot{\varepsilon},\dot{\zeta}]\|_{\mathfrak{h},\kappa,A}\big{)}+$
$\mathcal{O}_{A}L^{-j}|\varepsilon-\dot{\varepsilon}|\big{(}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}[\varepsilon,\zeta],K_{j}[\varepsilon,\dot{\zeta}],K_{j}[\dot{\varepsilon},\zeta],K_{j}[\dot{\varepsilon},\zeta]\|_{2\mathfrak{h},2\kappa,A}\big{)}.$ (50)
Both bounds still hold in the case $\varepsilon=\dot{\varepsilon}$ and $j=-\mathsf{N}-1$.
Regarding the last claim, we just recall that $K_{-\mathsf{N}-1}=0$ so that no integrability issues can arise. Anyway, note the nice structure of the second estimate: no term involves a difference of $K_{j}$ with both $\varepsilon$ and $\dot{\varepsilon}$ appearing. This enables us to bootstrap the estimates, as we first assume that $\varepsilon=\dot{\varepsilon}$, get some bounds, and then reinsert the bounds into the same estimate but with $\varepsilon\neq\dot{\varepsilon}$.
Corollary 2.4.
Suppose $L\gtrsim_{\sigma^{-1},\delta^{-1}}1$ (with other constants selected accordingly). We have $K_{j}\in\mathbb{B}^{j,2-\delta}_{2\mathfrak{h}_{*},2\kappa_{*},A}$ for all $j\geqslant-\mathsf{N}-1$, more precisely
$\|K_{j}\|_{2\mathfrak{h}_{*},2\kappa_{*},A}\leqslant\|\zeta_{0;j}\|^{2-\delta}_{\wp},$ (51)
and furthermore for $j\geqslant-\mathsf{N}$,
$\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h}_{*},\kappa_{*},A}\leqslant\|\zeta_{0;j},\dot{\zeta}_{0;j}\|^{1-\delta}_{\wp}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}L^{-j}|\varepsilon-\dot{\varepsilon}|\big{)}.$ (52)
Proof.
The first claim is verified inductively. For $j+1$,
$\|{K_{j+1}}\|_{2\mathfrak{h}_{*},2\kappa_{*},A}\leqslant\big{(}\mathcal{O}_{1}L^{2\sigma}+\mathcal{O}_{A}\|\zeta_{0;j}\|^{\delta}_{\wp}\big{)}\|\zeta_{0;j}\|^{2-\delta}_{\wp,(LX)^{[4]}}\leqslant\|\zeta_{0,j+1}\|^{2-\delta}_{\wp}$ (53)
as desired. For the second part, we first assume $\varepsilon=\dot{\varepsilon}$. Then
$\|K_{j+1}[\varepsilon,\zeta]$- $K_{j+1}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h}_{*},\kappa_{*},A}\leqslant$
$\mathcal{O}_{1}L^{2\sigma}\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h}_{*},\kappa_{*},A}+$
$\mathcal{O}_{A}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h}_{*},\kappa_{*},A}\big{)}.$ (54)
Inductively, in much the same way as above,
$\|K_{j+1}[\varepsilon,\zeta]-K_{j+1}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h}_{*},\kappa_{*},A}\leqslant\|\zeta_{0,j+1},\dot{\zeta}_{0;j+1}\|^{1-\delta}_{\wp}\|\zeta_{0;j+1}-\dot{\zeta}_{0;j+1}\|_{\wp}.$ (55)
Using this, we can bootstrap the estimates and, essentially repeating the arguments, find the result. ∎
This already looks like a Lipschitz property. Finishing up is not too complicated because there are only finitely many RG steps remaining. The next section deals with those.
2.4 Remaining steps
Recall that
$\mathrm{Z}(J)=e^{\frac{1}{2}\langle\mathrm{C}_{\varepsilon}J,J\rangle+E}\,\mathbb{E}\bigg{[}\mathcal{Z}_{j_{\bullet}}\!\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\mathrm{C}_{\varepsilon}J\bigg{)}\bigg{]}.$ (56)
We define ${\mathcal{Z}}^{\mathrm{red}}_{j}[\varepsilon,\zeta]=e^{-E[\varepsilon,\zeta]-\mathcal{E}_{j}[\varepsilon,\zeta]}\mathcal{Z}_{j}[\varepsilon,\zeta]$. Then
$\mathrm{Z}(J)=e^{\frac{1}{2}\langle\mathrm{C}_{\varepsilon}J,J\rangle}\frac{\mathbb{E}\big{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\big{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\mathrm{C}_{\varepsilon}J\big{)}\big{]}}{\mathbb{E}\big{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\big{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}\big{)}\big{]}}.$ (57)
The estimates we have should allow us to get the Lipschitz property, with the only catch being that we are dividing by a quantity that we also have to be able to bound away from 0. This is not trivial because $K_{j}$ may in general become negative, though of course their sum is still non-negative (even positive before taking the limit). But we first deal with the Gaussian prefactor and the numerator.
So, first compute
$\big{|}e^{\frac{1}{2}\langle\mathrm{C}_{\varepsilon}J,J\rangle}-e^{\frac{1}{2}\langle\mathrm{C}_{\dot{\varepsilon}}J,J\rangle}\big{|}$ $\leqslant|\langle(\mathrm{C}_{\dot{\varepsilon}}-\mathrm{C}_{\varepsilon})J,J\rangle|\int_{0}^{1}e^{\frac{1}{2}\langle((1-t)\mathrm{C}_{\varepsilon}+t\mathrm{C}_{\dot{\varepsilon}})J,J\rangle}\mathrm{d}t$
$\lesssim_{1}|\varepsilon-\dot{\varepsilon}|e^{\mathcal{O}_{1}\|J\|_{*}^{2}}.$ (58)
For the numerator,
$\|{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}-\dot{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\|_{\mathfrak{h}_{*};\langle\phi\rangle}$ $\leqslant\sum_{X\in\mathcal{P}_{j_{\bullet}}}\|e^{V_{j_{\bullet}}}-e^{\dot{V}_{j_{\bullet}}}\|_{{\mathfrak{h}_{*};\langle X^{c},\phi\rangle}}\prod_{Y\in\mathcal{C}(X)}\|K_{j_{\bullet}}\|_{{\mathfrak{h}_{*};\langle Y,\phi\rangle}}+$
$\hskip 15.0pt\sum_{X\in\mathcal{P}_{j_{\bullet}}}\|e^{\dot{V}_{j_{\bullet}}}\|_{{\mathfrak{h}_{*};\langle X^{c},\phi\rangle}}\prod_{Y\in\mathcal{C}(X)}\|\dot{K}_{j_{\bullet}}-K_{j_{\bullet}}\|_{{\mathfrak{h}_{*};\langle Y,\phi\rangle}}$
$\lesssim_{1}4^{|\mathbb{T}^{2}_{j_{\bullet}}|}G_{\kappa_{*}}\!(\mathbb{T}^{2}_{j_{\bullet}},\phi)\big{(}\|\zeta-\dot{\zeta}\|_{\wp}+|\varepsilon-\dot{\varepsilon}|\big{)}.$ (59)
Then
$\bigg{\|}$ $\mathbb{E}\bigg{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}(\cdot)\bigg{)}-\dot{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\dot{\varepsilon};k}+\mathrm{S}_{L^{-j_{\bullet}}}(\cdot)\bigg{)}\bigg{]}\bigg{\|}_{\mathfrak{h}_{*}/2;\langle\phi\rangle}$
$\leqslant\mathbb{E}\Big{[}\|{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}-\dot{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\|_{\mathfrak{h}_{*}/2;\big{\langle}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\phi\big{\rangle}}\Big{]}+$
$\hskip 20.0pt\mathbb{E}\bigg{[}\sup_{t\in[0,1]}\bigg{\|}\bigg{\langle}\mathrm{D}|_{\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}((1-t)\Psi_{\hskip-1.0pt\varepsilon;k}+t\Psi_{\hskip-1.0pt\dot{\varepsilon};k})+\mathrm{S}_{L^{-j_{\bullet}}}(\cdot)}\dot{\mathcal{Z}}_{j_{\bullet}}^{\mathrm{red}},\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}(\Psi_{\hskip-1.0pt\dot{\varepsilon};k}-\Psi_{\hskip-1.0pt\varepsilon;k})\bigg{\rangle}\bigg{\|}_{\mathfrak{h}_{*}/2;\langle\phi\rangle}\bigg{]}$
$\lesssim_{1}\mathcal{O}_{1}^{|\mathbb{T}^{2}_{j_{\bullet}}|}\big{(}\|\zeta-\dot{\zeta}\|_{\wp}+|\varepsilon-\dot{\varepsilon}|\big{)}\mathbb{E}\bigg{[}G_{\kappa_{*}}\!\bigg{(}\mathbb{T}^{2}_{j_{\bullet}},\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\phi\bigg{)}\bigg{]}+$
$\hskip 20.0pt\mathcal{O}_{1}^{|\mathbb{T}^{2}_{j_{\bullet}}|}\sum_{k=j_{\bullet}}^{0}\mathbb{E}\big{[}\|\Psi_{\hskip-1.0pt\dot{\varepsilon};k}-\Psi_{\hskip-1.0pt\varepsilon;k}\|_{\mathcal{C}^{2}(\mathbb{T}^{2}_{k})}^{2}\big{]}^{1/2}\cdot$
$\hskip 25.0pt\mathbb{E}\bigg{[}G_{\kappa_{*}/8}\!\bigg{(}\mathbb{T}^{2}_{j_{\bullet}},\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\phi\bigg{)}\bigg{]}^{\!1/4}\mathbb{E}\bigg{[}G_{\kappa_{*}/8}\!\bigg{(}\mathbb{T}^{2}_{j_{\bullet}},\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\dot{\varepsilon};k}+\mathrm{S}_{L^{-j_{\bullet}}}\phi\bigg{)}\bigg{]}^{\!1/4}$
$\leqslant\mathcal{O}_{1}^{|\mathbb{T}^{2}_{j_{\bullet}}|}G_{\kappa_{*}/8}(\mathbb{T}^{2},\phi)\big{(}\|\zeta-\dot{\zeta}\|_{\wp}+|\varepsilon-\dot{\varepsilon}|\big{)}.$ (60)
Finally, if we allow $\phi$ to vary,
$\bigg{|}\mathbb{E}\bigg{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}$ $\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\mathrm{C}_{\varepsilon}J\bigg{)}\bigg{]}-\mathbb{E}\bigg{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}\mathrm{C}_{\dot{\varepsilon}}J\bigg{)}\bigg{]}\bigg{|}$
$\leqslant\int_{0}^{1}\bigg{|}\bigg{\langle}\mathrm{D}|_{((1-t)\mathrm{C}_{\varepsilon}+\mathrm{C}_{\dot{\varepsilon}}t)J}\mathbb{E}\bigg{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}+\mathrm{S}_{L^{-j_{\bullet}}}(\cdot)\bigg{)}\bigg{]},(\mathrm{C}_{\dot{\varepsilon}}-\mathrm{C}_{\varepsilon})J\bigg{\rangle}\bigg{|}\mathrm{d}t$
$\leqslant\mathcal{O}_{1}^{|\mathbb{T}^{2}_{j_{\bullet}}|}|\varepsilon-\dot{\varepsilon}|e^{\mathcal{O}_{1}\|J\|^{2}_{*}}.$ (61)
Suppose now that
$m_{\bullet}\coloneqq\inf_{\varepsilon\leqslant\varepsilon_{\bullet},\zeta\in\mathrm{B}_{\bullet}}\mathbb{E}\bigg{[}{\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}\bigg{(}\sum_{k=j_{\bullet}}^{0}\mathrm{S}_{L^{k-j_{\bullet}}}\Psi_{\hskip-1.0pt\varepsilon;k}\bigg{)}\bigg{]}>0.$ (62)
This is the object of the next section. It then immediately follows that
$\|\mathrm{Z}[\varepsilon,\zeta]-\mathrm{Z}[\dot{\varepsilon},\dot{\zeta}]\|_{\mathfrak{C}}\leqslant\mathcal{O}_{1}^{|\mathbb{T}^{2}_{j_{\bullet}}|}m_{\bullet}^{-2}\big{(}\|\zeta-\dot{\zeta}\|_{\wp}+|\varepsilon-\dot{\varepsilon}|\big{)},$ (63)
which is exactly the Lipschitz property.
Let’s also consider here the small coupling case, i.e. $\zeta_{\bullet}=\zeta_{*}$, i.e. $j_{\bullet}=0$. Then
$\mathcal{Z}_{0}^{\mathrm{red}}(\phi)=e^{V_{0}(\mathbb{T}^{2},\phi)}+K_{0}(\mathbb{T}^{2},\phi).$ (64)
Doing a standard computation gives
$\mathrm{Z}[0,\zeta](J)=e^{\frac{1}{2}\langle\mathrm{C}J,J\rangle+V(\mathrm{C}J)}\frac{1+K^{V}_{1}\!(\mathrm{C}J)}{1+K^{V}_{1}\!(0)}$ (65)
where $V(\phi)=\int\zeta\big{(}1-\!\cos\!\big{(}\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\phi\big{)}\big{)}$ and
$\|K_{1}^{V}\|_{\mathfrak{h}_{*},\kappa_{*}/2}\lesssim_{1}\|\zeta\|_{\wp}^{2-\delta}.$ (66)
To show non-Gaussianity for small constant coupling, we can consider the connected truncated four-point function. This equals
$\mathfrak{S}^{\mathrm{c},\mathrm{t}}_{4}(f_{1},f_{2},f_{3},f_{4})$ $=\frac{\partial^{4}}{\partial t_{1}\partial t_{2}\partial t_{3}\partial t_{4}}\bigg{|}_{t_{1}=t_{2}=t_{3}=t_{4}=0}\big{(}V+\log(1+K^{V}_{1})\big{)}\bigg{(}\sum_{i=1}^{4}t_{i}f_{i}\bigg{)}$
$=-\beta^{2}\zeta\int_{\mathbb{T}^{2}}(f_{1}f_{2}f_{3}f_{4})(x)\mathrm{d}x+\sum_{\pi\in\mathrm{part}(4)}\frac{(-1)^{|\pi|-1}}{(1+K^{V}_{1}\!(0))^{|\pi|}}\prod_{I\in\pi}\bigg{\langle}\mathrm{D}^{|I|}|_{0}K^{V}_{1},\bigotimes_{i\in I}f_{i}\bigg{\rangle}$ (67)
[…]
$|\mathfrak{S}^{\mathrm{c},\mathrm{t}}_{4}(1,1,1,1)|\geqslant\beta^{2}|\zeta|-\mathcal{O}_{1}|\zeta|^{2-\delta}>0.$ (68)
3 The lover bound
Showing $m_{\bullet}>0$ requires some special theory, which we have sequestered in this section. Suppose we manage to find a neighborhood $U_{\bullet}$ of 0 in $\mathcal{C}^{\infty}(\mathbb{T}_{j_{\bullet}}^{2})$ such that
${\mathcal{Z}}^{\mathrm{red}}_{j_{\bullet}}|_{U_{\bullet}}\geqslant z_{\bullet}>0.$ (69)
It will then follow that
[…]
The probability appearing here is nonzero because a Gaussian measure assigns nonzero weight to every neighborhood of the origin, so $m_{\bullet}>0$. Hence we only need to find $U_{\bullet}$, which will of course be some kind of norm ball.
The idea is to apply polymer systems theory (cf. e.g. [GJ87], Chapter 20) to exponentiate the polymer representation. That is, if we define
$K_{j}^{V}\!(X,\phi)=e^{-V_{j}(X,\phi)}K_{j}(X,\phi),$ (71)
then
$\mathcal{Z}_{j}^{\mathrm{red}}(\phi)=\exp\!\bigg{(}V_{j}(\mathbb{T}^{2}_{j},\phi)+\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{(X_{1},\ldots,X_{n})\in\mathbf{C}_{n;j}}\mathfrak{n}(X_{1},\ldots,X_{n})K^{V}_{j}\!(X_{1},\phi)\ldots K^{V}_{j}\!(X_{n},\phi)\bigg{)}.$ (72)
[…]
The equality holds as long as the right-hand side converges. We cannot show this for all $\phi$ because $K_{j}^{V}$ could become large, so we define $U_{\bullet}$ by the condition $G_{\kappa_{*}}\!(\mathbb{T}^{2}_{j_{\bullet}},\phi)\leqslant 2$. Then
$\bigg{|}\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{(X_{1},\ldots,X_{n})\in\mathbf{C}_{n;j_{\bullet}}}$ $\mathfrak{n}(X_{1},\ldots,X_{n})K^{V}_{j_{\bullet}}\!(X_{1},\phi)\ldots K^{V}_{j_{\bullet}}\!(X_{n},\phi)\bigg{|}$
$\leqslant\sum_{n=1}^{\infty}\frac{\big{(}2\|\zeta_{j_{\bullet}}\|_{\wp}^{2-\delta}\big{)}^{\!n}}{n!}\sum_{(X_{1},\ldots,X_{n})\in\mathbf{C}_{n;j_{\bullet}},\tau\in\mathcal{T}_{n}}^{*}A^{-|X_{1}|}\ldots A^{-|X_{n}|}.$ (73)
[…]
$\sum_{n=1}^{\infty}$ $\frac{\big{(}2\|\zeta_{j_{\bullet}}\|_{\wp}^{2-\delta}\big{)}^{\!n}}{n!}\sum_{X_{1},\ldots,X_{n}\in\mathbf{C}_{n;j_{\bullet}},\tau\in\mathcal{T}_{n}}^{*}A^{-|X_{1}|}\ldots A^{-|X_{n}|}$
[…]
We start pruning the tree at the leaves, working inwards towards $i_{0}$. For instance, if $i$ is a vertex with leaves $j_{1},\ldots,j_{d_{i}-1}$ ($d_{i}$ being the degrees of $\tau$), then we first perform the sum ($B$ here is some arbitrary block in $X_{i}$ just serving to pin $X_{j_{1}}$)
$\sum_{X_{j_{1}}\in\mathcal{P}_{\mathrm{c};j_{\bullet}}}A^{-|X_{j_{1}}|}\mathbb{I}_{X_{i}\cap X_{j_{1}}\neq\varnothing}\leqslant|X_{i}|\sum_{X_{j_{1}}\in\mathcal{P}_{\mathrm{c};j_{\bullet}}}^{B\cap X_{j_{1}}\neq\varnothing}A^{-|X_{j_{1}}|}\leqslant 9|X_{i}|\sum_{k=1}^{\infty}(\mathcal{O}_{1}A^{-1})^{k}\leqslant|X_{i}|.$ (75)
After doing this for all the leaves, $i$ now becomes a leaf itself and a factor $|X_{i}|^{d_{i}-1}\leqslant(d_{i}-1)!\,e^{|X_{i}|}$ is produced. The exponential is merged with $A$ to give $(e^{-1}A)^{-|X_{i}|}$. These factors of $e^{-1}$ do not accumulate over time, since they are always merged to a different polymer. Anyway, proceeding like this (the last step when pruning the leaves of $i_{0}$ is slightly different but basically the same) leaves us with
$\sum_{n=1}^{\infty}\frac{\big{(}2\|\zeta_{j_{\bullet}}\|_{\wp}^{2-\delta}\big{)}^{\!n}}{n!}\sum_{X_{1},\ldots,X_{n}\in\mathbf{C}_{n;j_{\bullet}},\tau\in\mathcal{T}_{n}}^{*}$ $A^{-|X_{1}|}\ldots A^{-|X_{n}|}$
$\leqslant 2\zeta_{*}^{2-\delta}|\mathbb{T}^{2}_{j_{\bullet}}|\bigg{(}1+\sum_{n=2}^{\infty}\frac{\big{(}2\zeta_{*}^{2-\delta}\big{)}^{\!n-1}}{n!}\sum_{\tau\in\mathcal{T}_{n}}\prod_{i=1}^{n}(d_{i}-1)!\bigg{)}.$ (76)
Using Cayley’s formula for counting trees with degree sequence fixed and counting the number of possible degree sequences gives
$\sum_{\tau\in\mathcal{T}_{n}}\prod_{i=1}^{n}(d_{i}-1)!\leqslant 4^{n-1}(n-2)!$ (77)
and so we conclude that
$z_{\bullet}\geqslant e^{-2\zeta_{*}|\mathbb{T}^{2}_{j_{\bullet}}|}.$ (78)
Appendix A Scale decomposition
A.1 Continuous decomposition
Following [Bau13], we may realize the massive GFF on $\mathbb{T}^{2}$ as an integral against an $\mathrm{L}^{2}$-cylindrical Brownian motion $B_{t}$:
${\Phi}=\int_{0}^{\infty}t^{-1/2}{\mathrm{F}}_{\hskip-1.0ptt}(\cdot\,;\hskip-0.5pt1)*\mathrm{d}{B}_{t}.$ (79)
Here ${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5pt1)=0$ when $|x|_{2}\geqslant t/2$, and its precise form is
${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5ptm)=t\sum_{p\in 2\pi\mathbb{Z}}\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}$ (80)
where $\varphi$ is a real even $1/2$-Gevrey function [^5] We don’t need this much regularity, but we take it since it comes for free. with $\mathrm{supp}(\hat{\varphi})\subset[-1/2,1/2]$ normalized so that
$\int_{0}^{\infty}t\,\varphi(t)^{2}\mathrm{d}t=1.$ (81)
Let $\chi$ be the standard smooth bump function on $\mathbb{R}^{2}$ with support radius $1$. For $\varepsilon>0$, put $\chi_{\varepsilon}(x)=\varepsilon^{-2}\chi(\varepsilon^{-1}x)$. After mollifying and periodizing, we can decompose
$\Phi_{\hskip-0.5pt\varepsilon}$ $=\int t^{-1/2}{\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(\cdot\,;\hskip-0.5ptm)*\mathrm{d}{B}_{t},$
${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ $=(\chi_{\varepsilon}*{\mathrm{F}}_{\hskip-1.0ptt})(x)=t\hskip 1.0pt\sum_{p\in 2\pi\mathbb{T}^{2}}\hat{\chi}(\varepsilon p)\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}.$ (82)
The covariance decomposes as
$\mathrm{C}_{\varepsilon}(x)=(\chi_{\varepsilon}*\mathrm{C}*\chi_{\varepsilon})(x)=\int_{0}^{\infty}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)\frac{\mathrm{d}t}{t}.$ (83)
The following estimate is immediate for any multiindex $\alpha$ (it holds for $\varepsilon=0$ as well):
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}t^{-|\alpha|}e^{-\sqrt{tm}}.$ (84)
When $t<\varepsilon$, we expect better estimates since we have mollified. Indeed, in this case
$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}(\varepsilon^{-1}t)^{2}\varepsilon^{-|\alpha|}.$ (85)
Finally, we wish to know the value at $x=0$ more accurately since it enters directly into the RG flow. The sum over $2\pi\mathbb{Z}^{2}$ is a bit awkward, so we first work on all of $\mathbb{R}^{2}$: if $\bar{\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}$ is the analogue of ${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}$ on the full plane, straightforward computations with the Fourier transform show
$\bar{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(0;\hskip-0.5ptm)=\frac{1}{2\pi}+r^{(1)}_{\varepsilon;t}(m);\quad\quad|r^{(1)}_{\varepsilon;t}(m)|\lesssim_{1}\varepsilon t^{-1}+t^{2}m^{2}.$ (86)
Because of the Poisson summation formula
$\mathrm{F}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)=\sum_{y\in\mathbb{Z}^{2}}\bar{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x-y;\hskip-0.5ptm)$ (87)
and the finite range property, $\mathrm{F}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)=\bar{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ if $t<1/4$, so for these values the above result transfers immediately.
A.2 Discrete decomposition & scaling
Write $\varepsilon=\epsilon L^{-\mathsf{N}-1}/6$ for $\epsilon\in[1,L)$. Then put ($j=-\mathsf{N},\ldots,-1$)
$\mathrm{C}_{\varepsilon;-\mathsf{N-1}}(x)=\int_{0}^{L^{-\mathsf{N}}/6}\mathrm{F}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5pt1)\frac{\mathrm{d}t}{t},\quad\mathrm{C}_{\varepsilon;j}(x)=\int_{L^{j}/6}^{L^{j+1}/6}\mathrm{F}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5pt1)\frac{\mathrm{d}t}{t},\quad\mathrm{C}_{\varepsilon;0}(x)=\int_{1/6}^{\infty}\mathrm{F}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5pt1)\frac{\mathrm{d}t}{t}$ (88)
with corresponding fields $\Phi_{\hskip-0.5pt\varepsilon;j}$. The covariance $\mathrm{C}_{\varepsilon;j}$ for $j\leqslant-1$ has range at most $L^{j}/2$.
We will also rescale these covariances so that we are always working “at scale 11. Let $\mathrm{S}_{\lambda}$ be the scaling map $(\mathrm{S}_{\lambda}\phi)(x)=\phi(\lambda^{-1}x)$. Then
$\Gamma_{\hskip-1.0pt\varepsilon;j}=\mathrm{S}_{L^{-j}}\mathrm{C}_{\varepsilon;j}\mathrm{S}_{L^{-j}}^{*}$ (89)
with the adjoint taken in $\mathrm{L}^{2}$. The corresponding fields are denoted $\Psi_{\hskip-1.0pt\varepsilon;j}=\mathrm{S}_{L^{-j}}\Phi_{\hskip-0.5pt\varepsilon;j}$. It is easy to check that
$\Gamma_{\hskip-1.0pt\varepsilon;j}(x)=\mathrm{C}_{\hskip-1.0pt\varepsilon;j}(L^{j}x).$ (90)
If $j\leqslant-1$,
[…]
For the diagonal values $x=0$, we find that ($j\neq-\mathsf{N}-1$)
$\Gamma_{\hskip-1.0pt\varepsilon;j}(0)=\frac{\log(L)}{2\pi}+r^{(2)}_{\varepsilon;j};\quad\quad|r^{(2)}_{\varepsilon;j}|\lesssim_{1}1.$ (92)
This also implies that with $\zeta_{\varepsilon;j}$ as in (12),
$\zeta_{\varepsilon;j+1}=\varrho^{(1)}_{\varepsilon;j}L^{2\sigma}\mathrm{S}_{L^{-1}}\zeta_{\varepsilon;j}$ (93)
where $\varrho^{(1)}_{\varepsilon;j}$ is bounded uniformly away from $0,\infty$ if $j\neq-\mathsf{N}-1$. For the $j=0$ case, similar reasoning gives
$|\partial^{\alpha}\Gamma_{\hskip-1.0pt\varepsilon;0}(x)|\lesssim_{|\alpha|}1.$ (94)
Finally, we will also need to compare the fields at the same scales but with different covariances. Take $\dot{\varepsilon}\leqslant\varepsilon$ and restrict $j\neq-\mathsf{N}$. We have
$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j})|(x)$ $\lesssim_{|\alpha|}L^{-j}|\varepsilon-\dot{\varepsilon}|,$ (95)
$\mathbb{E}\big{[}\|\Psi_{\hskip-1.0pt\varepsilon;j}-\Psi_{\hskip-1.0pt\dot{\varepsilon};j}\|^{2}_{\mathcal{C}^{r}(X)}\big{]}$ $\lesssim_{r}L^{-2j}|\varepsilon-\dot{\varepsilon}|^{2}|X|,$ (96)
$\|\zeta_{\varepsilon;j}-\zeta_{\dot{\varepsilon};j}\|_{\wp}$ $\lesssim_{1}L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j}\|_{\wp},$ (97)
$\|\zeta_{\varepsilon;j}-\dot{\zeta}_{\varepsilon;j}\|_{\wp}$ $=\varrho^{(2)}_{\varepsilon;j}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}.$ (98)
Here $X\in\mathcal{P}_{j}$ and $\varrho^{(2)}_{\varepsilon;j}$ is bounded uniformly away from $0,\infty$. The last equality also holds for $j=-\mathsf{N}-1$, but then $\varrho^{(2)}_{\varepsilon;-\mathsf{N}-1}$ is bounded uniformly away from $0,\infty$ modulo a dependence on $L$. Showing all these relations is straightforward, using only Gaussian random variable theory, manipulations with the mollifier $\chi_{\varepsilon}$ in Fourier space, and the already stated properties of $\Gamma_{\hskip-1.0pt\varepsilon;j}$.
A.3 Finer scale decomposition
We will need a refined scale decomposition to prove some estimates below. Assume $L=\ell^{\mathfrak{m}}$ with $\mathfrak{m}\geqslant\ell^{2}$. As above, we have covariances (defined for $j=-\mathsf{N},\ldots,-1$, $s=-\mathfrak{m},\ldots,-1$):
$\mathrm{C}_{\varepsilon;j;s}(x)=\int_{L^{j+1}\ell^{s}/6}^{L^{j+1}\ell^{s+1}/6}\mathrm{F}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5pt1)\frac{\mathrm{d}t}{t}$ (99)
with attendant rescaled versions $\Gamma_{\hskip-1.0pt\varepsilon;j;s}(x)=\mathrm{C}_{\varepsilon;j;s}(L^{j+1}\ell^{s}x)$ and fields $\Psi_{\hskip-1.0pt\varepsilon;j;s}$. These again satisfy
[…]
$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j;s}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j;s})|(x)$ $\lesssim_{1}L^{-(j+1)}\ell^{-s}|\varepsilon-\dot{\varepsilon}|,$ (103)
$\mathbb{E}\big{[}\|\Psi_{\hskip-1.0pt\varepsilon;j}-\Psi_{\hskip-1.0pt\dot{\varepsilon};j}\|^{2}_{\mathcal{C}^{r}(X)}\big{]}$ $\lesssim_{r}L^{-2(j+1)}\ell^{-2s}|\varepsilon-\dot{\varepsilon}|^{2}|X|.$ (104)
Appendix B Polymers
B.1 Basic definitions
Let $\mathcal{B}_{j;s}$ denote all closed squares of the tesselation $\mathbb{Z}^{2}/L^{-j-1}\ell^{-s}\mathbb{Z}$ of $\mathbb{T}^{2}_{j;s}=L^{-j-1}\ell^{-s}\mathbb{T}^{2}$. $\mathcal{P}_{j;s}$ is the set of polymers which are finite unions (including $\varnothing$) of elements of $\mathcal{B}_{j;s}$. For $X\in\mathcal{P}_{j;s}$, we write $|X|$ for the area of $X$. $\mathcal{P}_{\mathrm{c};j;s}\subset\mathcal{P}_{j;s}$ denotes nonempty polymers which are connected as sets. $\mathcal{S}_{j;s}\subset\mathcal{P}_{\mathrm{c};j;s}$ denotes “small” polymers which are defined as being connected with $|X|\leqslant 4$. Collars of $X$, denoted $X^{[r]}$, are the sets of points which are at an $\infty$-distance of at most $r$ from $X$. $X^{*}=X^{[3]}$ is especially important as it equals the unions of all the small polymers intersecting $\mathrm{Int}(X)$. $\bar{X}^{(k)}$ denotes the smallest member of $\mathcal{P}_{j;s+k}$ (of course $s$ will “carry into” $j$ if $s+k\geqslant 0$) containing $\ell^{-k}X$. We use the special notation $\bar{X}^{L}=\bar{X}^{(\mathfrak{m})}$. $\mathcal{C}(X)$ denotes the connected components of $X$. Finally, we will most often deal with polymers in $\mathbb{T}^{2}_{j;-\mathfrak{m}}$, so in this case we drop the index $\mathfrak{m}$.
B.2 Geometric estimates
A general insight of the Wilsonian RG is that the effective potential stays roughly local. That is, we begin with a potential which is given by a local expression in the field and the RG flow is supposed to approximately preserve this property; this is why we have introduced the distinction between small and large polymers. Note, for instance, that if $X$ is not small, then $|\bar{X}^{(1)}|<|X|$, at least if $\ell$ is large enough. In fact, this inequality always gains us a factor uniform in $|X|$, which enables us to completely control the behavior on large sets and which is the content of the next lemma [SS09].
Lemma B.1.
If $X\in\mathcal{P}_{\mathrm{c};j;s}$ is not small and $k\geqslant 1$, then with $\eta=\frac{1}{101}$,
$|\bar{X}^{(k)}|\leqslant\frac{1}{1+\eta}|X|.$ (105)
Proof.
It’s enough to consider $k=1$. For clarity, we put $2=d$ (the proof works in all dimensions with $\eta$ depending on $d$, but we won’t bother thinking through the geometry there). We will prove (105) inductively with $\eta=(2^{d}+1+2^{d}\Delta)^{-1}$ where $\Delta=6^{d}-4^{d}+3^{d}-2^{d}-1$ is the largest number of blocks that can touch a connected polymer of $2^{d}+1$ blocks.
The induction will be on $|\bar{X}^{(1)}|$, starting with $|\bar{X}^{(1)}|=2^{d}+1$. Because $\ell$ is large, the only way we can have $|X|=|\bar{X}^{(1)}|$ is if $X$ is either disconnected or small, so we are forced to conclude that in our case, $|X|\geqslant 2^{d}+2$. Hence
$\bigg{(}1+\frac{1}{2^{d}+1}\bigg{)}|\bar{X}^{(1)}|\leqslant|X|,$ (106)
which witnesses the inductive base case. Now suppose $|\bar{X}^{(1)}|=n+1$. Remove from $\bar{X}$ a connected subset $Y$ comprising $2^{d}+1$ blocks. Then $\bar{X}-Y$ breaks up into at most $\Delta$ connected components (since every connected component certainly intersects the $1$-block thickening of $\bar{X}$), denoted $Z_{1},\ldots,Z_{\Delta}$ and ordered by decreasing size (we pad with $\varnothing$ if necessary). Let $k$ be the index such that $Z_{i}\geqslant 2^{d}+1$ when $i\leqslant k$ and $Z_{i}\leqslant 2^{d}$ for $i>k$. Let $M=\sum_{i=1}^{k}|Z_{i}|$ and $m=\sum_{i=k+1}^{\Delta}|Z_{i}|$. Then
$\frac{|X|}{|\bar{X}^{(1)}|}\geqslant\frac{2^{d}+2+(1+\eta)M+m}{2^{d}+1+M+m}\geqslant 1+\frac{1+\eta M}{2^{d}+1+M+2^{d}\Delta}=1+\eta.$ (107)
Due to the way we set up the RG, disconnected polymers sometimes appear in the formulas. To control these, we also have the following proposition [Fal12].
Lemma B.2.
Let $X$ be any polymer. Then
[…]
Proof.
Break $X$ up into $Y$ the small and $Z$ the large components. Since $|\bar{X}^{(j)}|\leqslant|\bar{Y}^{(j)}|+|\bar{Z}^{(j)}|$, we must have either $\frac{1}{2}|\bar{X}^{(j)}|\leqslant|\bar{Y}^{(j)}|$ or $\frac{1}{2}|\bar{X}^{(j)}|\leqslant|\bar{Z}^{(j)}|$. In the former case we simply get
$|\bar{X}^{(j)}|\leqslant 2|\bar{Y}^{(j)}|\leqslant 2|Y|\leqslant 8|\mathcal{C}(X)|,$ (109)
while in the latter we apply the previous lemma:
$|X|=|Y|+|Z|\geqslant|\bar{Y}^{(j)}|+(1+\eta)|\bar{Z}^{(j)}|\geqslant|\bar{X}^{(j)}|+\eta|\bar{Z}^{(j)}|\geqslant(1+\eta/2)|\bar{X}^{(j)}|.$ (110)
B.3 $\mathcal{C}^{r}$ spaces & regulators
The basic function norms for $X\in\mathcal{P}_{j;s}$ are
$\|\nabla^{n}\hskip-1.0ptf\|_{\mathrm{L}^{p}(X)}$ $=\big{\|}(\|\partial^{\mu}\!f\|_{\mathrm{L}^{p}(X)})_{|\mu|=n}\big{\|}_{\ell^{p}},$ (111)
$\|\nabla^{n}\hskip-1.0ptf\|_{\mathrm{L}^{p}(\partial X)}$ $=\big{\|}(\|\partial^{\mu}\!f\|_{\mathrm{L}^{p}(\partial X)})_{|\mu|=n}\big{\|}_{\ell^{p}},$ (112)
$\|\nabla^{n}\hskip-1.0ptf\|_{\mathcal{B}^{q}_{p}(X)}$ $=\big{\|}(\|\nabla^{n}\hskip-1.0ptf\|_{\mathrm{L}^{p}(B)})_{B\in\mathcal{B}_{j}(X)}\big{\|}_{\ell^{q}},$ (113)
$\|\nabla^{n}\hskip-1.0ptf\|_{\mathcal{C}^{r}(X)}$ $=\max_{0\leqslant m\leqslant r}\|\nabla^{n+m}\hskip-1.0ptf\|_{\mathrm{L}^{\infty}(X)},$ (114)
$\|f\|_{\wp}$ $=\|f\|_{\mathcal{B}^{\wp}_{\infty}(\mathbb{T}^{2}_{j})}.$ (115)
The last norm will only be used to control the coupling; note that when $j=0$ it reduces to the $\mathrm{L}^{\wp}$ norm. We can also see that $\|\!\cdot\!\|_{\mathcal{B}^{p}_{p}}=\|\!\cdot\!\|_{\mathrm{L}^{p}}$.
The notion of differentiable function on $X$ requires a bit of care since this is generically not a submanifold with corners of $\mathbb{T}^{2}$. However, it is the closure of a relatively compact open set and it is rather easy to convince oneself that it is a Whitney domain, [^6] The only problematic points could be where two squares touch at only one vertex, but the angle at which they meet is $90^{\circ}\neq 0$, so the Whitney condition holds. This is also true of polymers in higher dimensions. so the extension theorem applies (for instance see [Whi34], though this only treats the case of subsets in $\mathbb{R}^{2}$; the extension theorem for the torus can easily be obtained by periodization arguments). For instance, all reasonable definitions of $\mathcal{C}^{r}(X)$ coincide and have good properties: we may define this space as either those functions in $\mathcal{C}^{r}(\mathrm{Int}(X))$ whose $r$ derivatives all have continuous extensions to the boundary, or equivalently as those functions on $X$ which are restrictions of $\mathcal{C}^{r}$ functions on $\mathbb{T}^{2}$.
We will also need functionals of the field values. To control their growth in the field, we introduce regulators [SS09, Fal12] depending on $\kappa>0$ (for $\phi\in\mathcal{C}^{2}(X)$)
$\kappa\log G_{\kappa}(X,\phi)$ $=\mathfrak{c}^{(1)}\big{(}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla^{2}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}\big{)}+\mathfrak{c}^{(2)}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial X)},$ (116)
$\kappa\log g_{\kappa}(X,\phi)$ $=\mathfrak{c}^{(3)}\sum_{n=0}^{2}\|\nabla^{n}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}.$ (117)
$G_{\kappa}$ is what appears directly in norms, while $g_{\kappa}$ is an important intermediate quantity in computations.
The actual functionals are smooth functions $F$ on $\mathcal{C}^{2}(X)$, but for now we consider a general Banach space $\mathcal{X}$. Then if $\mathfrak{h}>0$ and $x\in\mathcal{X}$, we put
$\|F\|_{n;\langle x\rangle;\mathcal{X}}$ $=\sup_{y_{1},\ldots,y_{n}\in\mathrm{Ball}(\mathcal{X})}\big{|}\langle\mathrm{D}^{n}|_{x}F,y_{1},\ldots,y_{n}\rangle\big{|},$ (118)
$\|F\|_{\mathfrak{h};\langle x\rangle;\mathcal{X}}$ $=\sum_{n=0}^{\infty}\frac{\mathfrak{h}^{n}}{n!}\|F\|_{n;\langle x\rangle;\mathcal{X}}.$ (119)
We also denote by $\mathcal{N}_{\mathfrak{h}}(\mathcal{X})$ the set of functionals for which $\|F\|_{\mathfrak{h};\langle x\rangle}<\infty$ for all $x$.
Finally, the norms satisfy some identities that we will use without further comment. If $T:\mathcal{X}\rightarrow\mathcal{Y}$ is bounded,
$\|F\circ T\|_{\mathfrak{h};\langle x\rangle;\mathcal{X}}\leqslant\|F\|_{\|T\|\mathfrak{h};\langle Tx\rangle;\mathcal{Y}}.$ (120)
For instance, $\mathrm{S}_{\lambda}:\mathcal{C}^{r}(X)\rightarrow\mathcal{C}^{r}(\lambda X)$ has norm $1$ if $\lambda\geqslant 1$. Furthermore, if $q_{1,2}:\mathcal{X}\rightarrow\mathcal{Y}_{1,2}$ are two quotients and $F_{1,2}\in\mathcal{N}_{\mathfrak{h}}(\mathcal{Y}_{1,2})$,
$\|(F_{1}\circ q_{1})(F_{2}\circ q_{2})\|_{\mathfrak{h};\langle\phi\rangle;\mathcal{X}}\leqslant\|F_{1}\|_{\|q_{1}\|\mathfrak{h};\langle q_{1}\phi\rangle;\mathcal{Y}_{1}}\|F_{2}\|_{\|q_{2}\|\mathfrak{h};\langle q_{2}\phi\rangle;\mathcal{Y}_{2}}.$ (121)
For instance, $\mathcal{C}^{r}(X\cup Y)$ quotients with norm $1$ to $\mathcal{C}^{r}(X)$ and $\mathcal{C}^{r}(Y)$, so we have a kind of product inequality for functionals on $\mathcal{C}^{r}$-spaces where the underlying sets can change.
B.4 Polymer activities
[…]
$\|K\|_{n;\langle X,x\rangle;\mathcal{X}_{X}}$ $=\|K(X)\|_{n;\langle x\rangle;\mathcal{X}_{X}},$ (122)
$\|K\|_{\mathfrak{h};\langle X,x\rangle;\mathcal{X}_{X}}$ $=\|K(X)\|_{\mathfrak{h};\langle x\rangle;\mathcal{X}_{X}},$ (123)
$\|K\|_{\mathfrak{h},w;\langle X\rangle;\mathcal{X}_{X}}$ $=\sup_{x\in\mathcal{X}_{X}}w_{X}(x)^{-1}\|K\|_{\mathfrak{h};\langle X,x\rangle;\mathcal{X}_{X}},,$ (124)
$\|K\|_{\mathfrak{h},w,A;\mathcal{X}_{X}}$ $=\sup_{X\in\mathcal{P}_{j}}A^{|X|}\|K\|_{\mathfrak{h},w;\langle X\rangle;\mathcal{X}_{X}}.$ (125)
The subscript $\mathcal{X}_{X}$ on the left-hand sides is retained only if we feel the need to emphasize what the underlying Banach spaces are, but we will always drop it when $\mathcal{X}_{X}=\mathcal{C}^{2}(X)$ which is most often the case. If $w$ depends on a further parameter (for us, $w_{X}=G_{\kappa}(X)$), we write that parameter in place of $w$ in (124), (125). If $\alpha>0$, write $\mathbb{B}^{j,\alpha}_{\mathfrak{h},w,A}$ for the set of polymer activities which satisfy
$\|K\|_{\mathfrak{h},w,A;\mathcal{X}_{X}}\leqslant\|(\zeta_{\bullet})_{0,j}\|^{\alpha}_{\wp}=\Big{(}L^{2j}e^{\frac{1}{2}\beta\sum_{k=j}^{0}\Gamma_{\hskip-1.0pt0;j}}\zeta_{\bullet}\Big{)}^{\!\alpha}.$ (126)
B.4.1 Complexification of functionals
In one proof, we will need a “complexification” map to get contractive estimates on charged terms. The whole development of this section is rather elementary, comprising mostly Taylor expansions and basic multilinear algebra.
[…]
$^{\mathbb{C\hskip-1.0pt}}F(w)=\sum_{n=0}^{\infty}\frac{1}{n!}\langle\mathrm{D}^{n}|_{x}F,z,\ldots,z\rangle$ (127)
where $\mathrm{D}^{n}|_{x}F$ is of course extended to complex arguments by multilinearity. The series converges: if $z=z_{0}+iz_{1}$ are the real and imaginary parts,
[…]
$\leqslant\mathfrak{h}^{n}\|F\|_{n;\langle X\rangle}.$ (128)
We still need to check that this is independent of the decomposition of $w$, so suppose $w=x+z=x^{\prime}+z^{\prime}$. Then $\|x-x^{\prime}\|_{\mathcal{X}}<\mathfrak{h}$ and by Taylor expansion,
$\langle\mathrm{D}^{n}|_{x}F,u_{1},\ldots,u_{n}\rangle=\sum_{m=0}^{\infty}\frac{1}{m!}\langle\mathrm{D}^{n+m}|_{x^{\prime}}F,x-x^{\prime},\ldots,x-x^{\prime},u_{1},\ldots,u_{n}\rangle.$ (129)
Hence
$^{\mathbb{C\hskip-1.0pt}}F(x+z)$ $=\sum_{n,m=0}^{\infty}\frac{1}{n!m!}\langle\mathrm{D}^{n+m}|_{x^{\prime}}F,x-x^{\prime},\ldots,x-x^{\prime},z,\ldots,z\rangle$
$=\sum_{k=0}^{\infty}\frac{1}{k!}\sum_{n=0}^{k}\binom{k}{n}\langle\mathrm{D}^{k}|_{x^{\prime}}F,x-x^{\prime},\ldots,x-x^{\prime},z,\ldots,z\rangle$
$=\sum_{k=0}^{\infty}\frac{1}{k!}\langle\mathrm{D}^{k}|_{x^{\prime}}F,x-x^{\prime}+z,\ldots,x-x^{\prime}+z\rangle$
$={{}^{\mathbb{C\hskip-1.0pt}}F}(x^{\prime}+z^{\prime}).$ (130)
This shows well-definedness.
Some further properties: if $w$ is already real, $^{\mathbb{C\hskip-1.0pt}}F(w)=F(w)$. $^{\mathbb{C\hskip-1.0pt}}F$ is also a smooth functional on $\mathcal{X}^{\mathbb{C},\mathfrak{h}}$: if $u_{1},\ldots,u_{n}\in\mathcal{X}^{\mathbb{C}}$,
$\frac{\partial^{m}}{\partial t_{1}\ldots\partial t_{m}}\bigg{|}_{t_{1}=\ldots=t_{m}=0}$ ${{}^{\mathbb{C\hskip-1.0pt}}F}\bigg{(}w+\sum_{i=1}^{m}t_{i}u_{i}\bigg{)}=$
$=$ $\frac{\partial^{m}}{\partial t_{1}\ldots\partial t_{m}}\bigg{|}_{t_{1}=\ldots=t_{m}=0}\sum_{n=0}^{\infty}\frac{1}{n!}\bigg{\langle}\mathrm{D}^{n}|_{x}F,\psi+\sum_{i=1}^{m}t_{i}u_{i},\ldots,\psi+\sum_{i=1}^{m}t_{i}u_{i}\bigg{\rangle}$
$=$ $\sum_{n=0}^{\infty}\frac{1}{n!}\langle\mathrm{D}^{n+m}|_{x}F,\psi,\ldots,\psi,u_{1},\ldots,u_{m}\rangle$ (131)
as we could have guessed from (129). Furthermore, with $z$ pure imaginary we can derive
$\big{\|}{{}^{\mathbb{C\hskip-1.0pt}}F}\big{\|}_{m;\langle w\rangle}\leqslant 2^{m}\sum_{n=0}^{\infty}\frac{\|z\|^{n}_{\mathcal{X}}}{n!}\|F\|_{n+m;\langle x\rangle}$ (132)
and so
$\big{\|}{{}^{\mathbb{C\hskip-1.0pt}}F}\big{\|}_{\mathfrak{k};\langle w\rangle}\leqslant\|F\|_{2\mathfrak{k}+\|\mathrm{Im}(w)\|_{\mathcal{X}};\langle\mathrm{Re}(w)\rangle},$ (133)
which is finite as long as $2\mathfrak{k}+\|\mathrm{Im}(w)\|_{\mathcal{X}}\leqslant\mathfrak{h}$.
Appendix C Norm & regulator estimates
C.1 Algebraic inequalities
The next proposition comprises a menagerie of norm inequalities of various flavors which serve as the technical basis of the paper. This is all standard [SS09]; (145) is probably the most interesting: it allows us to absorb multiplicative factors into norms with improved scaling compared to plain $\mathrm{L}^{2}(X)$ which ultimately enables us to prove (159), a strong estimate used to kill polynomials in the field that arise when Taylor expanding polymer activities as in (169).
Lemma C.1.
Let $f,g\in\mathcal{C}^{2}(\mathbb{T}^{2})$ and $X,Y\in\mathcal{P}_{j;s},B\in\mathcal{B}_{j;s}$, $\alpha>0$. $\mathrm{diam}(X)$ is taken relative to paths entirely within $X$. The following inequalities hold:
$\|f\|_{\mathrm{L}^{\infty}(X)}^{2}$ $\lesssim_{1}\sum_{n=0}^{2}\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathrm{L}^{2}(X)},$ (134)
$\|f\|_{\mathrm{L}^{2}(X)},\|f\|_{\mathrm{L}^{2}(X)}$ $\lesssim_{1}\|f\|_{\mathcal{B}^{2}_{\infty}(X)},$ (135)
$Y\subset X\in\mathcal{P}_{\mathrm{c};j;s}:\|f\|^{2}_{\mathrm{L}^{\infty}(X)}$ $\leqslant\frac{2}{|Y|}\|f\|^{2}_{\mathrm{L}^{2}(Y)}+2\,\mathrm{diam}(X)^{2}\|\nabla^{2}\hskip-1.0ptf\|_{\mathcal{B}^{2}_{\infty}(X)}^{2},$ (136)
$X\in\mathcal{S}:\|\nabla f\|^{2}_{\mathcal{C}^{1}(X)}$ $\lesssim_{1}\|\nabla f\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathcal{B}^{2}_{\infty}(X)},$ (137)
$\|f\|_{\mathrm{L}^{2}(\partial B)}^{2}$ $\lesssim_{1}\|f\|^{2}_{\mathrm{L}^{2}(B)}+\|\nabla f\|^{2}_{\mathrm{L}^{\infty}(B)},$ (138)
$\|f\|^{2}_{\mathrm{L}^{2}(\partial X)}$ $\leqslant\|f\|^{2}_{\mathrm{L}^{2}(\partial\ell\bar{X}^{(1)})}+\mathcal{O}_{1}\big{(}\|f\|^{2}_{\mathrm{L}^{2}(\ell\bar{X}^{(1)}-X)}+\|\nabla f\|^{2}_{\mathcal{B}^{2}_{\infty}(\ell\bar{X}^{(1)}-X)}\big{)},$ (139)
$\|\nabla(f+g)\|^{2}_{\mathrm{L}^{2}(X)}$ $\lesssim\|\nabla f\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}\alpha\big{(}\|\nabla f\|^{2}_{\mathrm{L}^{2}(\partial X)}+\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathrm{L}^{2}(X)}\big{)}$
$\hskip 15.0pt+\|\nabla g\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}\alpha^{-1}\big{(}\|g\|^{2}_{\mathrm{L}^{2}(\partial X)}+\|g\|^{2}_{\mathrm{L}^{2}(X)}\big{)}.$ (140)
Furthermore, we have the scaling relations
$\|\nabla^{n}\mathrm{S}_{\ell}f\|_{\mathcal{B}^{q}_{p}(\ell X)}$ $\leqslant\ell^{\frac{2}{p\wedge q}-n}\|\nabla^{n}\hskip-1.0ptf\|_{\mathcal{B}^{q}_{p}(X)},$ (141)
$\|\nabla^{n}\mathrm{S}_{\ell}f\|_{\mathrm{L}^{p}(\partial\ell X)}$ $=\ell^{\frac{1}{p}-n}\|\nabla^{n}\hskip-1.0ptf\|_{\mathrm{L}^{p}(\partial X)},$ (142)
$\|\mathrm{S}_{\ell}f\|_{\wp}$ $\leqslant\ell^{2/\wp}\|f\|_{\wp},$ (143)
$\|\mathrm{S}_{\ell^{-1}}f\|_{\wp}$ $\leqslant\ell^{2/\wp}\|f\|_{\wp}.$ (144)
Analogous estimates hold with $L$ in place of $\ell$.
Finally, take $X=X_{1}$ small and fix a $\lambda>1$. If $\ell>\mathcal{O}_{\lambda}$, there are disjoint translates $X_{2},\ldots,X_{32\lceil\lambda\rceil}\subset\ell\bar{X}^{(1)}$ of $X_{1}$ with $\mathfrak{X}=\bigsqcup_{i}X_{i}$ such that
$\lambda\|f\|_{\mathrm{L}^{2}(X)}^{2}\leqslant\|f\|^{2}_{\mathrm{L}^{2}(\mathfrak{X})}+\mathcal{O}_{\lambda}\|\nabla f\|^{2}_{\mathcal{B}^{2}_{\infty}(\ell\bar{X}^{(1)})}.$ (145)
Proof.
The scaling relations are all straightforward. Now consider a block $B\in\mathcal{B}_{j}$. We may assume it lies in the first quadrant and touches the origin $(0,0)$. Define $h(x)=x_{1}x_{2}f(x)$ so that
$h(x)=\int_{0}^{x_{2}}\int_{0}^{x_{1}}\partial_{1}\partial_{2}h(y_{1},y_{2})\mathrm{d}y_{1}\mathrm{d}y_{2}$ (146)
and
$x_{1}x_{2}|f(x)|\leqslant\|\partial_{1}\partial_{2}h\|_{\mathrm{L}^{1}(B)}\leqslant\sum_{n=0}^{2}\|\nabla^{2}\hskip-1.0ptf\|_{\mathrm{L}^{1}(B)}.$ (147)
if $x\in[1/2,1]^{2}$ we divide by the factor $x_{1}x_{2}$, otherwise we appeal to symmetry. So by Hölder,
$\|f\|^{2}_{\mathrm{L}^{\infty}(B)}\leqslant 3\sum_{n=0}^{2}\|\nabla^{n}\hskip-1.0ptf\|^{2}_{\mathrm{L}^{2}(B)},$ (148)
proving (134) after summing over $B\in\mathcal{B}_{j}(X)$. For (135), we simply compute
$\|f\|^{2}_{\mathrm{L}^{2}(\partial X)}\leqslant\sum_{B\in\mathcal{B}_{j}(X)}\|f\|_{\mathrm{L}^{2}(\partial B)}\leqslant 4\sum_{B\in\mathcal{B}_{j}(X)}\|{f}\|_{\mathrm{L}^{\infty}(B)}^{2}=4\|f\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}$ (149)
and analogously for $\|f\|_{\mathrm{L}^{2}(X)}$. Next, it is obvious that for $x,y\in X$, $X$ connected,
$|f(x)|\leqslant|f(y)|+\mathrm{diam}(X)\|\nabla f\|_{\mathrm{L}^{\infty}(X)}.$ (150)
If $y\in Y$, by squaring and taking the average over $y$, we get (136). Alternatively, fixing a point $x$ in the boundary of a block $B$ and a point $y\in B$,
$|f(x)|^{2}\lesssim_{1}|f(y)|^{2}+\|\nabla f\|^{2}_{\mathrm{L}^{\infty}(B)}$ (151)
and again averaging, first over $y$ and then over $x\in\partial B$, we obtain (138). For (137), by (136),
$\|\nabla f\|^{2}_{\mathcal{C}^{1}(X)}\leqslant\|\nabla f\|^{2}_{\mathrm{L}^{\infty}(X)}+\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathrm{L}^{\infty}(X)}\lesssim_{1}\|\nabla f\|_{\mathrm{L}^{2}(X)}^{2}+\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}.$ (152)
Now observe that (139) follows from (138),
$\|f\|^{2}_{\mathrm{L}^{2}(\partial X)}$ $\leqslant\|f\|^{2}_{\mathrm{L}^{2}(\partial\ell\bar{X}^{(1)})}+\|f\|^{2}_{\mathrm{L}^{2}(\partial(\ell\bar{X}^{(1)}-X))}$
$\leqslant\|f\|^{2}_{\mathrm{L}^{2}(\partial\ell\bar{X}^{(1)})}+\mathcal{O}_{1}\big{(}\|f\|^{2}_{\mathrm{L}^{2}(\ell\bar{X}^{(1)}-X)}+\|\nabla f\|^{2}_{\mathcal{B}^{2}_{\infty}(\ell\bar{X}^{(1)}-X)}\big{)}.$ (153)
For (140), we apply integration by parts to find
$\|\nabla(f+g)\|^{2}_{\mathrm{L}^{2}(X)}$ $=\|\nabla f\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla g\|^{2}_{\mathrm{L}^{2}(X)}+2\sum_{|\mu|=1}\langle\partial^{\mu}\!f,\partial^{\mu}\!g\rangle_{\mathrm{L}^{2}(X)}$
$\leqslant\|\nabla f\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla g\|^{2}_{\mathrm{L}^{2}(X)}+2\sum_{|\mu|=1}\!\big{(}|\langle\alpha\partial^{2\mu}\!f,\alpha^{-1}g\rangle_{\mathrm{L}^{2}(X)}|+|\langle\alpha\partial^{\mu}\!f,\alpha^{-1}g\rangle_{\mathrm{L}^{2}(\partial X)}|\big{)}.$ (154)
By Cauchy-Schwarz and the arithmetic-geometric inequality,
$\|\nabla(f+g)\|^{2}_{\mathrm{L}^{2}(X)}$ $\leqslant\|\nabla f\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla g\|^{2}_{\mathrm{L}^{2}(X)}$
$\hskip 15.0pt+4\alpha\big{(}\|\nabla^{2}\hskip-1.0ptf\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla f\|^{2}_{\mathrm{L}^{2}(\partial X)}\big{)}+8\alpha^{-1}\big{(}\|g\|^{2}_{\mathrm{L}^{2}(X)}+\|g\|^{2}_{\mathrm{L}^{2}(\partial X)}\big{)}.$ (155)
For (145), we first have to find the polymer translates. This is easy: as long as $\ell\geqslant\mathcal{O}_{\lambda}$, we can just pick some sidelength-$\ell$ square of $\ell\bar{X}^{(1)}$ which $X$ intersects and make copies of $X$ inside that square. By keeping the copies close together, the radius of the resulting pattern will be of order $\lambda$ (even $\sqrt{\lambda}$ since we are in 2d). Then by (136),
$\lambda\|f\|^{2}_{\mathrm{L}^{2}(X)}$ $\leqslant\|f\|^{2}_{\mathrm{L}^{2}(X_{1})}+\frac{\lambda-1}{32\lceil\lambda\rceil-1}\sum_{k=2}^{32\lceil\lambda\rceil}\|f\|^{2}_{\mathrm{L}^{2}(X_{1})}$
$\leqslant\|f\|^{2}_{\mathrm{L}^{2}(X_{1})}+\frac{|X_{1}|^{2}}{32}\sum_{k=2}^{32\lceil\lambda\rceil}\|f\|^{2}_{\mathrm{L}^{\infty}(X_{1})}$
$\leqslant\|f\|^{2}_{\mathrm{L}^{2}(X_{1})}+\sum_{k=2}^{32\lceil\lambda\rceil}\|f\|^{2}_{\mathrm{L}^{2}(X_{k})}+\mathcal{O}_{\lambda}\|f\|^{2}_{\mathrm{L}^{\infty}(\ell\bar{X}^{(1)})}$
$\leqslant\|f\|^{2}_{\mathrm{L}^{2}(\mathfrak{X})}+\mathcal{O}_{\lambda}\|\nabla f\|^{2}_{\mathcal{B}^{2}_{\infty}(\ell\bar{X}^{(1)})},$ (156)
concluding the proof. ∎
As stated, we will use (145) in the next proposition to remove an otherwise problematic constant. In that case $\lambda=2$ so all the $\mathcal{O}_{\lambda}\!$’s become $\mathcal{O}_{1}\!$’s. The following estimates on regulators are still of a purely algebraic nature, but are necessary to have good control of the fluctuation step.
Lemma C.2.
Because $\mathfrak{c}^{(2)}\lesssim_{1}1\lesssim_{1}\mathfrak{c}^{(1)}\wedge\mathfrak{c}^{(3)}\wedge\ell$, we have the estimates
$X\subset Y:G_{\kappa}(X,\phi)$ $\leqslant G_{\kappa}(Y,\phi),$ (157)
$X\in\mathcal{P}_{\mathrm{c};j;s}:G_{\kappa}(X,\mathrm{S}_{\ell}\phi+\psi)$ $\leqslant G_{\kappa}(\bar{X}^{(1)},\phi)g_{\kappa}(X,\psi),$ (158)
$X\in\mathcal{S}_{j;s}:G_{\kappa}(X,\mathrm{S}_{\ell}\phi+\psi)$ $\leqslant G_{\kappa}(\bar{X}^{(1)},\phi)^{1/2}g_{\kappa}(X,\psi).$ (159)
Proof.
We can take $\kappa=1$. For (157), by induction take $Y=X\cup B$ with $B$ a single square. Since
$\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}\leqslant\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial Y)}+\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial B)},$ (160)
by (138),
$\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}\leqslant\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial Y)}+\mathcal{O}_{1}\big{(}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(B)}+\|\nabla^{2}\phi\|^{2}_{\mathrm{L}^{\infty}(B)}\big{)},$ (161)
so by (136),
$\log G_{1}(X,\phi)$ $\leqslant\mathfrak{c}^{(1)}\big{(}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(X)}+\|\nabla^{2}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}\big{)}+\mathfrak{c}^{(2)}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial Y)}+\mathcal{O}_{1}\mathfrak{c}^{(2)}\big{(}\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(B)}+\|\nabla^{2}\phi\|^{2}_{\mathrm{L}^{\infty}(B)}\big{)}$
[…]
$\leqslant\log G_{1}(Y,\phi)$ (162)
where we have had to require $\mathfrak{c}^{(1)}\gtrsim_{1}\mathfrak{c}^{(2)}$. Moving on, by (135,140) and the triangle inequality,
$\log G_{1}(X,\mathrm{S}_{\ell}\phi+\psi)\leqslant\,$ $\mathfrak{c}^{(1)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}\alpha\mathfrak{c}^{(1)}\big{(}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}+\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}\big{)}+$
$\mathfrak{c}^{(1)}\|\nabla\psi\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}\alpha^{-1}\mathfrak{c}^{(1)}\big{(}\|\psi\|^{2}_{\mathrm{L}^{2}(\partial X)}+\|\psi\|^{2}_{\mathrm{L}^{2}(X)}\big{)}+$
$2\mathfrak{c}^{(1)}\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}+2\mathfrak{c}^{(1)}\|\nabla^{2}\psi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}+2\mathfrak{c}^{(2)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}+2\mathfrak{c}^{(2)}\|\nabla\psi\|^{2}_{\mathrm{L}^{2}(\partial X)}$
$\leqslant\,$ $\mathfrak{c}^{(1)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}+(\mathcal{O}_{1}\alpha+2)\mathfrak{c}^{(1)}\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}+\mathcal{O}_{1}(\alpha\mathfrak{c}^{(1)}+\mathfrak{c}^{(2)})\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}+$
[…]
We estimate the third term by (139):
$\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}\leqslant\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial\ell\bar{X}^{(1)})}+\mathcal{O}_{1}\big{(}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\ell\bar{X}^{(1)}-X)}+\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(\ell\bar{X}^{(1)}-1)}\big{)}.$ (164)
Plugging this in and rescaling,
[…]
$\ell^{-1}\mathcal{O}_{1}(\alpha\mathfrak{c}^{(1)}+\mathfrak{c}^{(2)})\|\nabla\phi\|^{2}_{\mathrm{L}^{2}(\partial\bar{X}^{(1)})}+$
[…]
Now if $X$ is small and the conditions on constants as above, take (163) and combine it with (139,145):
$2\log G_{1}(X,\mathrm{S}_{\ell}\phi+\psi)\leqslant\,$ $2\mathfrak{c}^{(1)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}(\alpha+1)\mathfrak{c}^{(1)}\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}+$
[…]
To restore the constants, proceed as before. ∎
C.2 Fluctuation step bounds
We combine these estimates from the previous section with properties of the scale decomposition. These are a mix of [Fal12] and [DH00]; because it considers the critical phase, the former also does an analysis of the second order, which is much more complicated than the first-order analysis we have to perform.
Lemma C.3.
Take $X$ connected and let $\Psi_{\hskip-1.0pt\varepsilon;j},\Psi_{\hskip-1.0pt\varepsilon;j;s}$ be the various fluctuation fields. First take $j\leqslant-1$.
$\mathbb{E}\big{[}g_{\kappa_{*}/8}(X,\Psi_{\varepsilon;j;s})\big{]}\leqslant 2^{\log_{L}(\ell)|X|}.$ (167)
Furthermore, let $\kappa\geqslant\kappa_{*}/2$. Then
$\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa}(X,\mathrm{S}_{L}\phi+(1-t)\Psi_{\hskip-1.0pt\varepsilon;j}+\Psi_{\hskip-1.0pt\dot{\varepsilon};j})\Big{]}$ $\leqslant 2^{|X|}G_{\kappa}(\bar{X}^{L},\phi),$ (168)
$X\in\mathcal{S}:\big{(}1+\|\nabla\phi\|_{\mathcal{C}^{1}(\bar{X}^{L})}\big{)}^{\!2\,}\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa}(X,t\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\Big{]}$ $\lesssim_{1}\kappa G_{\kappa}(\bar{X}^{L},\phi).$ (169)
In case $j=0$, if $\kappa\geqslant\kappa_{*}/8$,
$\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa}(X,\phi+(1-t)\Psi_{\hskip-1.0pt\varepsilon;0}+\Psi_{\hskip-1.0pt\dot{\varepsilon};0})\Big{]}\leqslant 2G_{\kappa/2}(\mathbb{T}^{2},\phi).$ (170)
We note that (168) immediately implies the weaker
$\mathbb{E}\big{[}G_{\kappa}(X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\leqslant 2^{|X|}G_{\kappa}(\bar{X}^{L},\phi).$ (171)
Proof.
For (167), first take $|\mu|\leqslant 4$ and estimate
$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)}\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})).$ (172)
This holds as long as the $\mathrm{L}^{2}$ operator norm of $\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})$ is bounded by $1/2$, that is, we have to take $\kappa_{*}$ such that $\mathfrak{c}^{(3)}\|\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})\|_{\mathcal{B}(\mathrm{L}^{2}(X))}\lesssim_{1}\kappa_{*}$. Since
$\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})(f)(x)=\int_{\mathbb{T}_{j;s}^{2}}\mathbb{I}_{x,y\in X}(-\partial^{2\mu}\Gamma_{\hskip-1.0pt\varepsilon;j;s})(x-y)f(y)\mathrm{d}y,$ (173)
we apply Young’s inequality and the finite-range property to find
$\|\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})\|_{\mathcal{B}(\mathrm{L}^{2}(X))}\leqslant\|\partial^{2\mu}\Gamma_{\hskip-1.0pt\varepsilon;j;s}\|_{\mathrm{L}^{1}(\mathbb{T}^{2})}\lesssim_{1}\ell^{2}\|\partial^{2\mu}\Gamma_{\hskip-1.0pt\varepsilon;j;s}\|_{\mathrm{L}^{\infty}(\mathbb{T}^{2})}\lesssim_{1}\mathfrak{m}\log(\ell)=\log(L).$ (174)
Hence we need $\kappa_{*}\gtrsim_{1}\mathfrak{c}^{(3)}\log(L)$ which we do have since we took $\mathfrak{c}^{(4)}\gtrsim_{1}\mathfrak{c}^{(3)}$.
Now, because $\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})=|X|(-\partial^{2\mu}\Gamma_{\hskip-1.0pt\varepsilon;j;s})(0)$,
$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)}\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\log(\ell)|X|).$ (175)
By (134),
$\log g_{\kappa_{*}/8}(X,\Psi_{\hskip-1.0pt\varepsilon;j;s})\lesssim_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\sum_{n=0}^{4}\sum_{B\in\mathcal{B}_{j}(X)}\|\nabla^{n}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(B)}\lesssim_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\sum_{|\mu|\leqslant 4}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)}.$ (176)
There are $31$ terms in this sum, so by Hölder and $\mathfrak{c}^{(4)}\gtrsim_{1}\mathfrak{c}^{(3)}$, we get (167):
$\mathbb{E}\big{[}g_{\kappa_{*}/8}(X,\Psi_{\hskip-1.0pt\varepsilon;j;s})\big{]}\leqslant\prod_{|\mu|\leqslant 4}\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)})\big{]}^{\hskip-1.0pt1/31}\leqslant 2^{\log_{L}(\ell)|X|}.$ (177)
For (168) we argue inductively:
$\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa}(X,\mathrm{S}_{L}\phi,(1-t)\Psi_{\hskip-1.0pt\varepsilon;j}+\Psi_{\hskip-1.0pt\dot{\varepsilon};j})\Big{]}$ $=\mathbb{E}\bigg{[}\sup_{t\in[0,1]}G_{\kappa}\bigg{(}X,\mathrm{S}_{L}\phi+\sum_{s=0}^{\mathfrak{m}-1}\mathrm{S}_{\ell^{s}}\big{(}(1-t)\Psi_{\hskip-1.0pt\varepsilon;j;s}+t\Psi_{\hskip-1.0pt\dot{\varepsilon};j;s}\big{)}\bigg{)}\bigg{]}$
$\leqslant G_{\kappa}(\bar{X}^{L},\phi)\prod_{s=0}^{\mathfrak{m}-1}\mathbb{E}\Big{[}\sup_{t\in[0,1]}g_{\kappa}(\bar{X}^{(s)},(1-t)\Psi_{\hskip-1.0pt\varepsilon;j;s}+t\Psi_{\hskip-1.0pt\dot{\varepsilon};j;s})\Big{]}$
$\leqslant G_{\kappa}(\bar{X}^{L},\phi)\prod_{s=0}^{\mathfrak{m}-1}\mathbb{E}\big{[}g_{\kappa_{*}/8}(\bar{X}^{(s)},\Psi_{\hskip-1.0pt\varepsilon;j;s})\big{]}^{1/2}\mathbb{E}\big{[}g_{\kappa_{*}/8}(\bar{X}^{(s)},\Psi_{\hskip-1.0pt\dot{\varepsilon};j;s})\big{]}^{1/2}$
$\leqslant 2^{\log_{L}(\ell)\sum_{s=0}^{\mathfrak{m}-1}|\bar{X}^{(s)}|}G_{\kappa}(\bar{X}^{L},\phi)$
$\leqslant 2^{\log_{L}(\ell)\mathfrak{m}|X|}G_{\kappa}(\bar{X}^{L},\phi)$
$=2^{|X|}G_{\kappa}(\bar{X}^{L},\phi).$ (178)
For (169), first note that ($\kappa\geqslant 1$)
$(1+x)^{a}\lesssim_{a}\kappa^{a/2}e^{\kappa^{-1}x^{2}/2},$ (179)
so (taking $\mathfrak{c}^{(1)}\gtrsim_{1}1$ since we wish to apply (137))
$\big{(}1+\|\nabla\phi\|_{\mathcal{C}^{1}(\bar{X}^{L})}\big{)}^{\!2\,}\lesssim_{1}\kappa G_{\kappa}(\bar{X}^{L},\phi)^{1/2}.$ (180)
Proceeding inductively as before but applying (159) at the last step,
$\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa}(X,t\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\Big{]}$ $\leqslant\prod_{s=0}^{\mathfrak{m}-1}\mathbb{E}\big{[}g_{\kappa_{*}/2}(\bar{X}^{(s)},\Psi_{\hskip-1.0pt\varepsilon;j;s})\big{]}\sup_{t\in[0,1]}G_{\kappa}(\bar{X}^{L},t\phi)^{1/2}$
$\lesssim_{1}G_{\kappa}(\bar{X}^{L},t\phi)^{1/2}.$ (181)
By (180), the proof is finished, except the $j=0$ case. Here similar but simpler reasoning works (since $\kappa$ is of order $\log(L)$ and the last covariance is of order $\mathcal{O}_{1}$, there are no issues). ∎
Having established the basic estimates, we can obtain the crucial scaling properties of the polymer activities. The first proposition is very general and so very suboptimal; we have to use more model-specific ideas to beat the scaling factors coming up here.
Lemma C.4.
Let $X\in\mathcal{P}_{\mathrm{c};j}$, $\kappa\geqslant\kappa_{*}/2$. Then
$\big{\|}\mathbb{E}\big{[}K_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h},\kappa;\langle X\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\leqslant 2^{|X|}\|K_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}.$ (182)
Proof.
We just unravel the definitions.
$\big{|}\mathrm{D}^{n}|_{\phi}\mathbb{E}\big{[}K_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{|}$ $=\Bigg{|}\frac{\partial^{n}}{\partial t_{1}\ldots\partial t_{n}}\bigg{|}_{t_{1}=\ldots=t_{n}=0}\mathbb{E}\bigg{[}K_{j}\bigg{(}X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j}+\sum_{i=1}^{n}t_{i}\mathrm{S}_{L}f_{i}\bigg{)}\bigg{]}\Bigg{|}$
$\leqslant\mathbb{E}\big{[}\big{|}\langle\mathrm{D}^{n}|_{\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j}}K_{j}(X),\mathrm{S}_{L}f_{1},\ldots,\mathrm{S}_{L}f_{n}\rangle\big{|}\big{]}$
$\leqslant\mathbb{E}\big{[}\|K\|_{n;\langle X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j}\rangle}\big{]}\|f_{1}\|_{\mathcal{C}^{2}(\bar{X}^{L})}\ldots\|f_{n}\|_{\mathcal{C}^{2}(\bar{X}^{L})}.$ (183)
So $\big{\|}\mathbb{E}\big{[}K_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\leqslant\mathbb{E}\big{[}\|K\|_{\mathfrak{h};\langle X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j}\rangle}\big{]}$. By introducing $G_{\kappa}$ and using (171),
$\big{\|}\mathbb{E}\big{[}K_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\leqslant\mathbb{E}\big{[}G_{\kappa}(X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\|K\|_{\mathfrak{h},\kappa;\langle X\rangle}\leqslant 2^{|X|}G_{\kappa}(\bar{X}^{L},\phi)\|K\|_{\mathfrak{h},\kappa;\langle X\rangle}.$ (184)
The stronger scaling, as we’ve said, requires model-specific ideas. We will control small sets by using a strong contraction property for charged terms and by normalizing neutral terms by extracting the vacuum energy contribution. In the following propositions, $\mathrm{Rem}$ denotes the Taylor remainder.
Lemma C.5.
Let $X\in\mathcal{P}_{\mathrm{c};j}$.
$\|\mathrm{Rem}_{m}K_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\big{(}1+\mathfrak{h}^{-1}\|\phi\|_{\mathcal{C}^{2}(X)}\big{)}^{\!m+1\,}\sup_{t\in[0,1]}\sum_{n=m+1}^{\infty}\frac{\mathfrak{h}^{n}}{n!}\|K_{j}\|_{n;\langle X,t\phi\rangle}.$ (185)
Furthermore if $X\in\mathcal{S}$ and $x_{0}\in X$, denote $\delta\phi(x)=\phi(x)-\phi(x_{0})$. Then
$\|K(\cdot,\mathrm{S}_{L}\delta(\cdot)+\psi)\|_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\leqslant\|K\|_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi+\psi\rangle}.$ (186)
Proof.
For (186), we can take $x_{0}\in X$ which has a path joining it to any other point of lenght at most $2\sqrt{2}$. If $f\in\mathcal{C}^{2}(X)$,
[…]
$\leqslant L^{-1}\mathrm{diam}(X)\max_{m=1,2}\|\mathrm{S}_{L}\nabla^{m}\hskip-1.0ptf\|_{\mathrm{L}^{\infty}(X)}$
$\leqslant 8L^{-1}\|f\|_{\mathcal{C}^{2}(\bar{X}^{L})}.$ (187)
Therefore
$\big{|}\langle\mathrm{D}^{n}|_{\phi}K_{j}(X,\mathrm{S}_{L}\delta(\cdot)+\psi),f_{1},\ldots,f_{n}\rangle\big{|}$ $=\Bigg{|}\frac{\partial^{n}}{\partial t_{1}\ldots\partial t_{n}}\bigg{|}_{t_{1}=\ldots=t_{n}=0}K_{j}\bigg{(}X,\mathrm{S}_{L}\delta\phi+\psi+\sum_{i=1}^{n}t_{i}\mathrm{S}_{L}\delta f_{i}\bigg{)}\Bigg{|}$
$\leqslant(8L^{-1})^{n}\|K_{j}\|_{n;\langle X,\mathrm{S}_{L}\phi+\psi\rangle}\|f_{1}\|_{\mathcal{C}^{2}(\bar{X}^{L})}\ldots\|f_{n}\|_{\mathcal{C}^{2}(\bar{X}^{L})}.$ (188)
This indeed implies the desired inequality. For (185), we break up derivatives into two cases. For terms with $n\geqslant m+1$, $\mathrm{D}^{n}\mathrm{Rem}_{m}F=\mathrm{D}^{n}F$ since the remainder is obtained by subtracting a polynomial. On the other hand for $n\leqslant m$,
$\big{|}\langle\mathrm{D}^{n}|_{\phi}\mathrm{Rem}_{m}K_{j}(X),f_{1},\ldots,f_{n})\rangle\big{|}$ $\leqslant\int_{0}^{1}\frac{(1-t)^{m-n}}{(m-n)!}\big{|}\langle\mathrm{D}^{m+1}|_{t\phi}K_{j}(X),\phi,\ldots,\phi,f_{1},\ldots,f_{n}\rangle\big{|}\mathrm{d}t$
$\leqslant\frac{1}{(m+1-n)!}\sup_{t\in[0,1]}\|K_{j}\|_{m+1;\langle X,\phi\rangle}\|\phi\|_{\mathcal{C}^{2}(X)}^{m+1-n}\prod_{i=1}^{n}\|f_{i}\|_{\mathcal{C}^{2}(X)}.$ (189)
Then
$\sum_{n=0}^{m+1}\frac{\mathfrak{h}^{n}}{n!}\|\mathrm{Rem}_{m}K_{j}\|_{n;\langle X,\phi\rangle}$ $\leqslant\mathfrak{h}^{m+1}\sup_{t\in[0,1]}\|K_{j}\|_{m+1;\langle X,t\phi\rangle}\sum_{n=0}^{m+1}\frac{(\mathfrak{h}^{-1}\|\phi\|_{\mathcal{C}^{2}(X)})^{m+1-n}}{n!(m+1-n)!}$
$\leqslant\frac{\mathfrak{h}^{m+1}}{(m+1)!}\big{(}1+\mathfrak{h}^{-1}\|\phi\|_{\mathcal{C}^{2}(X)}\big{)}^{\!m+1\,}\sup_{t\in[0,1]}\|K_{j}\|_{m+1;\langle X,t\phi\rangle}$ (190)
and the result follows. ∎
Lemma C.6.
Let $|q|\geqslant 1$, $X\in\mathcal{S}$. Also fix $\mathfrak{h}_{*}\gtrsim_{1}1$ and take $\mathfrak{h}\geqslant\mathfrak{h}_{*}$. Then
$\big{\|}\mathbb{E}\big{[}k_{j}^{(q)}(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h},\kappa;\langle X\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}|q|\mathfrak{h}-(|q|-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}}2^{|X|}\|K_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}.$ (191)
Proof.
We only consider $q>0$. Take $f\in\mathcal{S}$ positive with integral $1$. We do an analytic continuation argument: if $z\in\mathbb{R}$ and $x_{0}\in X$,
$\mathbb{E}\big{[}k_{j}^{(q)}(X,\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$ $=\int k_{j}^{(q)}(X,\phi+\psi)\,\gamma[\Gamma_{\hskip-1.0pt\varepsilon;j}](\mathrm{d}\psi)$
$=e^{-\frac{1}{2}z^{2}\langle\Gamma_{\hskip-1.0pt\varepsilon;j}f,f\rangle}\int e^{-z\langle f,\psi\rangle}k_{j}^{(q)}(X,\phi+\psi+z\Gamma_{\hskip-1.0pt\varepsilon;j}f)\,\gamma[\Gamma_{\hskip-1.0pt\varepsilon;j}](\mathrm{d}\psi)$ (192)
$=e^{-\frac{1}{2}z^{2}\langle\Gamma_{\hskip-1.0pt\varepsilon;j}f,f\rangle+i\beta qz(\Gamma_{\hskip-1.0pt\varepsilon;j}f)(x_{0})}\int e^{-z\langle f,\psi\rangle}\,{{}^{\mathbb{C\hskip-1.0pt}}}k^{(q)}_{j}(X,\phi+\psi+z\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}f))\,\gamma[\Gamma_{\hskip-1.0pt\varepsilon;j}](\mathrm{d}\psi).$
By the identity principle, equality holds for all $z$ on some strip around $\mathbb{R}$. We note that
$\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\|\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}f)\|_{\mathcal{C}^{2}(X)}\lesssim_{1}\|\nabla\Gamma_{\hskip-1.0pt\varepsilon;j}f\|_{\mathcal{C}^{1}(X)}\lesssim_{1}\max_{|\mu|=1,2}\|\partial^{\mu}\Gamma_{\hskip-1.0pt\varepsilon;j}\|_{\mathrm{L}^{\infty}(\mathbb{T}^{2})}\lesssim_{1}1,$ (193)
so because $\mathfrak{h}_{*}\gtrsim_{1}1$, the strip is big enough that we can take $z=i\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}$. In the limit as $f$ approaches $\delta_{x_{0}}$,
$\mathbb{E}\big{[}k_{j}^{(q)}(X,\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}=e^{-(q-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\int e^{-i\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}\psi(x_{0})}\,{{}^{\mathbb{C\hskip-1.0pt}}}k^{(q)}_{j}(X,\phi+\psi+i\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}\delta_{x_{0}}))\,\gamma[\Gamma_{\hskip-1.0pt\varepsilon;j}](\mathrm{d}\psi).$ (194)
It follows that
$\big{\|}\mathbb{E}\big{[}k^{(q)}_{j}$ $(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}=\big{\|}e^{i\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathrm{S}_{L}(\cdot)(x_{0})}\mathbb{E}\big{[}k^{(q)}_{j}(\cdot,\mathrm{S}_{L}\delta(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}$
$\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathfrak{h}}\big{\|}\mathbb{E}\big{[}k_{j}^{(q)}(\cdot,\cdot+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi\rangle}$
$\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathfrak{h}-(q-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\big{\|}\mathbb{E}\big{[}e^{-i\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}\Psi_{\hskip-1.0pt\varepsilon;j}(x_{0})}\,{{}^{\mathbb{C\hskip-1.0pt}}}k^{(q)}_{j}(\cdot,\cdot+\Psi_{\hskip-1.0pt\varepsilon;j}+i\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}\delta_{x_{0}}))\big{]}\big{\|}_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi\rangle}$
$\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathfrak{h}-(q-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\mathbb{E}\big{[}\|{{}^{\mathbb{C\hskip-1.0pt}}}k^{(q)}_{j}\big{\|}_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi+\Psi_{\hskip-1.0pt\varepsilon;j}+i\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}\delta_{x_{0}})}\big{]}$
$\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathfrak{h}-(q-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\mathbb{E}\big{[}\|{{}^{\mathbb{C\hskip-1.0pt}}}k^{(q)}_{j}\big{\|}_{8\mathfrak{h}/L+2\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}\|\delta(\Gamma_{\hskip-1.0pt\varepsilon;j}\delta_{x_{0}})\|_{\mathcal{C}^{2}(X)};\langle X,\mathrm{S}_{L}\delta\phi+\Psi_{\hskip-1.0pt\varepsilon;j}\rangle}\big{]}$
$\leqslant e^{\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}q\mathfrak{h}-(q-1/2)\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\mathbb{E}\big{[}G_{\kappa}(X,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\|k^{(q)}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}$ (195)
where the last inequality holds due to $L,\mathfrak{h}_{*}\gtrsim_{1}1$. The proof concludes after applying (171). ∎
Lemma C.7.
Take $\kappa\geqslant\kappa_{*}/2$. Suppose $K_{j}$ is real and even in $\phi$. Then
$\big{\|}\mathrm{Rem}_{0}\mathbb{E}\big{[}k^{(0)}_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}\big{\|}_{\mathfrak{h},\kappa;\langle X\rangle;\mathcal{C}^{2}(\bar{X}^{L})}\lesssim_{1}2^{|X|}\kappa L^{-2}\|K\|_{\mathfrak{h},\kappa;\langle X\rangle}.$ (196)
Proof.
By assumption, $\mathrm{Rem}_{0}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}=\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$. From (185,186),
$\big{\|}\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\psi)\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}$ $\leqslant\|\mathrm{Rem}_{1}k^{(0)}_{j}\|_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi+\psi\rangle}$
$\leqslant\big{(}1+L\|\mathrm{S}_{L}\delta\phi\|_{\mathcal{C}^{2}(X)}\big{)}^{\!2\,}\sup_{t\in[0,1]}\sum_{n=2}^{\infty}\frac{(8\mathfrak{h}/L)^{n}}{n!}\|k^{(0)}_{j}\|_{n;\langle X,t\mathrm{S}_{L}\delta\phi+\psi\rangle}$
$\lesssim_{1}\big{(}1+\|\nabla\phi\|_{\mathcal{C}^{1}(X)}\big{)}^{\!2\,}L^{-2}\sup_{t\in[0,1]}\|k^{(0)}_{j}\|_{\mathfrak{h};\langle X,t\mathrm{S}_{L}\delta\phi+\psi\rangle}$ (197)
$\lesssim_{1}L^{-2}\|K_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}\big{(}1+\|\nabla\phi\|_{\mathcal{C}^{1}(X)}\big{)}^{\!2\,}\sup_{t\in[0,1]}G_{\kappa}(X,t\mathrm{S}_{L}\phi+\psi).$
The result is immediate by (169). ∎
Appendix D Proofs of some propositions
These proofs closely follow those in [Fal12]. In some respects they are simpler because we are in the subcritical regime, while in others they are more complicated because we have to deal with the fluctuation fields changing a bit at all scales when the UV cutoff is lowered.
D.1 Proof of Proposition 2.2
First,
$\langle\mathrm{D}^{n}|_{\phi}V_{j}(X),f_{1},\ldots,f_{n}\rangle=\beta^{n/2}\int_{X}\zeta_{j}(x)\cos^{(n)}\!\!\big{(}\sqrt{\hskip-1.16pt\beta\hskip 1.16pt}\phi(x)\big{)}f_{1}(x)\ldots f_{n}(x)\mathrm{d}x,$ (198)
so
$\|V_{j}-\dot{V}_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\lesssim_{1}|X|\|\zeta_{\varepsilon;j}-\dot{\zeta}_{\dot{\varepsilon},j}\|_{\wp}$
$\lesssim_{1}|X|\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}.$ (199)
Furthermore,
$\|e^{V_{j}}-e^{\dot{V}_{j}}\|_{\mathfrak{h};\langle X,\phi\rangle}$ $\leqslant\sum_{n=1}^{\infty}\frac{1}{n!}\|V_{j}^{n}-\dot{V}^{n}_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}$
$\leqslant\|V_{j}-\dot{V}_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}\sum_{n=1}^{\infty}\frac{1}{n!}\sum_{k=0}^{n-1}\|V_{j}\|^{n-1-k}_{\mathfrak{h};\langle X,\phi\rangle}\|\dot{V}_{j}\|^{k}_{\mathfrak{h};\langle X,\phi\rangle}$
$\leqslant e^{\|V_{j},\dot{V}_{j}\|^{n}_{\mathfrak{h};\langle X,\phi\rangle}}\|V_{j}-\dot{V}_{j}\|_{\mathfrak{h};\langle X,\phi\rangle}$
$\lesssim_{1}2^{|X|}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}.$ (200)
For the extraction,
$|Q_{j}$ $(X)-\dot{Q}_{j}(X)|$
$=\big{|}\mathbb{E}\big{[}k^{(0)}_{j}(X,\Psi_{\hskip-1.0ptj})-\dot{k}^{(0)}_{j}(X,\dot{\Psi}_{\hskip-1.0ptj})\big{]}\big{|}$
$\leqslant\mathbb{E}\big{[}\big{|}k^{(0)}_{j}-\dot{k}^{(0)}_{j}\big{|}(X,\Psi_{\hskip-1.0ptj})\big{]}+\mathbb{E}\big{[}\big{|}\dot{k}^{(0)}_{j}(X,\Psi_{\hskip-1.0ptj})-\dot{k}^{(0)}_{j}(X,\dot{\Psi}_{\hskip-1.0ptj})\big{|}\big{]}$
$\leqslant\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}\mathbb{E}\big{[}G_{\kappa}(X,\Psi_{\hskip-1.0ptj})\big{]}+\mathbb{E}\bigg{[}\int_{0}^{1}\big{|}\langle\mathrm{D}|_{(1-t){\Psi}_{\hskip-1.0ptj}+t\dot{\Psi}_{\hskip-1.0ptj}}\dot{k}^{(0)}_{j}(X),{\Psi}_{\hskip-1.0ptj}-\dot{\Psi}_{\hskip-1.0ptj}\rangle\big{|}\mathrm{d}t\bigg{]}$
$\lesssim_{1}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}+\|\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}\mathbb{E}\Big{[}\|{\Psi}_{\hskip-1.0ptj}-\dot{\Psi}_{\hskip-1.0ptj}\|_{\mathcal{C}^{2}(X)}\sup_{t\in[0,1]}G_{\kappa}(X,(1-t){\Psi}_{\hskip-1.0ptj}+t\dot{\Psi}_{\hskip-1.0ptj})\Big{]}$
$\leqslant\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}+\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}\mathbb{E}\big{[}\|{\Psi}_{\hskip-1.0ptj}-\dot{\Psi}_{\hskip-1.0ptj}\|_{\mathcal{C}^{2}(X)}\big{]}^{\!1/2}\mathbb{E}\Big{[}\sup_{t\in[0,1]}G_{\kappa/2}(X,(1-t){\Psi}_{\hskip-1.0ptj}+t\dot{\Psi}_{\hskip-1.0ptj})\Big{]}^{\!1/2}$
$\lesssim_{1}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa;\langle X\rangle}.$ (201)
This also immediately implies the estimate for $\delta\mathcal{E}_{j}$. For $J_{j}^{\mathfrak{B}\rightarrow X}-\dot{J}_{j}^{\mathfrak{B}\rightarrow X}$,
$|J_{j}^{\mathfrak{B}\rightarrow X}-\dot{J}_{j}^{\mathfrak{B}\rightarrow X}|\leqslant\sum_{\varnothing\neq\mathcal{A}\subset\mathcal{C}(X)}\prod_{Y\in\mathcal{A}}|J_{j}(Y,\mathfrak{B}_{Y})-\dot{J}_{j}(Y,\mathfrak{B}_{Y})|\prod_{Z\in\mathcal{A}^{c}}|\dot{J}_{j}(Z,\mathfrak{B}_{Z})|$ (202)
and so it will be enough to control $J_{j}(Y,B)-\dot{J}_{j}(Y,B)$, which is easily done from its definition:
$|J_{j}(Y,B)-\dot{J}_{j}(Y,B)|$ $\leqslant\bigg{(}\sum_{D\in\mathcal{B}_{j}(LB),X\in\mathcal{S}_{j}}^{D\subset X,\bar{X}=Y}+\,\mathbb{I}_{Y=B}\sum_{Y^{\prime}\in\mathcal{S}_{j+1}}\sum_{D\in\mathcal{B}_{j}(LB),X\in\mathcal{S}_{j}}^{D\subset X,\bar{X}=Y^{\prime}}\bigg{)}|Q_{j}(X)-\dot{Q}_{j}(X)|$
$\lesssim_{1}A^{-1}L^{2}\big{(}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}$
$\leqslant\mathcal{O}_{A}A^{-2|Y^{*}|}\big{(}\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}.$ (203)
In going to the last line, we used that $|Y^{*}|\leqslant 49|Y|<200$ because $Y\in\mathcal{S}$. Writing the inequality in this specific form comes in handy later.
Now let’s consider $R_{j}^{(\cdot)}$. Expanding the definition,
$\|R$ $^{(\cdot)}_{j}(\cdot,\psi)-\dot{R}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}$
$=\bigg{\|}\prod_{Y\in\mathcal{C}(X)}\bigg{(}\tilde{K}_{j}(Y,\cdot,\psi)-\sum_{B\in\mathcal{B}_{j+1}(Y)}J_{j}(Y,B)\bigg{)}-\prod_{Y\in\mathcal{C}(X)}\bigg{(}\dot{\tilde{K}}_{j}(Y,\cdot,\dot{\psi})-\sum_{B\in\mathcal{B}_{j+1}(Y)}\dot{J}_{j}(Y,B)\bigg{)}\bigg{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(X)}$
$\leqslant\sum_{\mathcal{A}\subset\mathcal{C}(X)}\bigg{\|}\prod_{Y\in\mathcal{A}}\tilde{K}_{j}(Y,\cdot,\psi)\sum_{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}J_{j}^{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}-\prod_{Y\in\mathcal{A}}\dot{\tilde{K}}_{j}(Y,\cdot,\dot{\psi})\sum_{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}\dot{J}_{j}^{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}\bigg{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(X)}$
$\leqslant\sum_{\varnothing\neq\mathcal{A}\subset\mathcal{C}(X)}\big{\|}\hat{K}_{j}^{\bigcup\mathcal{A}}(\cdot,\psi)-\dot{\hat{K}}_{j}^{\bigcup\mathcal{A}}(\cdot,\dot{\psi})\big{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(\bigcup\mathcal{A})}\sum_{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}|J_{j}^{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}|+$
$\hskip 45.0pt\sum_{\mathcal{A}\subsetneq\mathcal{C}(X)}\big{\|}\dot{\hat{K}}_{j}^{\bigcup\mathcal{A}}(\cdot,\dot{\psi})\big{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(\bigcup\mathcal{A})}\sum_{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}|J_{j}^{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}-\dot{J}_{j}^{\mathfrak{B}\rightarrow\bigcup\mathcal{A}^{c}}|.$ (204)
Here we have provisionally defined
$\hat{K}_{j}^{X}(\phi,\psi)=\prod_{Y\in\mathcal{C}(X)}\tilde{K}_{j}(Y,\phi,\psi).$ (205)
The main issue is controlling the norm of this object. We need to get a handle on $\tilde{K}_{j}$ itself (first with $\psi=\dot{\psi}$):
$\|\tilde{K}_{j}(\cdot$ $,\cdot,\psi)-\dot{\tilde{K}}_{j}(\cdot,\cdot,\psi)\|_{\mathfrak{h};\langle X,\phi\rangle}$
$\leqslant\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X}\bigg{\|}e^{V_{j}(LX-Y,\cdot)}\prod_{X^{\prime}\in\mathcal{C}(Y)}K_{j}(X^{\prime},\cdot)-e^{\dot{V}_{j}(LX-Y,\cdot)}\prod_{X^{\prime}\in\mathcal{C}(Y)}\dot{K}_{j}(X^{\prime},\cdot)\bigg{\|}_{\mathfrak{h};\langle\mathrm{S}_{L}\phi+\psi\rangle;\mathcal{C}^{2}(LX)}$
$\leqslant\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X}\|e^{V_{j}}-e^{\dot{V}_{j}}\|_{\mathfrak{h};\langle LX-Y,\mathrm{S}_{L}\phi+\psi\rangle}\prod_{X^{\prime}\in\mathcal{C}(Y)}\|K_{j}\|_{\mathfrak{h};\langle X^{\prime},\mathrm{S}_{L}\phi+\psi\rangle}+$
$\hskip 20.0pt\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X}\|e^{\dot{V}_{j}}\|_{\mathfrak{h};\langle LX-Y,\mathrm{S}_{L}\phi+\psi\rangle}\bigg{\|}\prod_{X^{\prime}\in\mathcal{C}(Y)}K_{j}(X^{\prime},\cdot)-\prod_{X^{\prime}\in\mathcal{C}(Y)}\dot{K}_{j}(X^{\prime},\cdot)\bigg{\|}_{\mathfrak{h};\langle\mathrm{S}_{L}\phi+\psi\rangle;\mathcal{C}^{2}(Y)}$
$\leqslant\mathcal{O}_{L}^{|X|}G_{\kappa}(LX,\mathrm{S}_{L}\phi+\psi)\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X}A^{-|Y|}\big{(}\mathcal{O}_{1}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}^{\!|\mathcal{C}(Y)|-1}\cdot$
$\hskip 40.0pt\big{(}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}$
$\leqslant\mathcal{O}_{A}A^{-(1+\eta/3)|X|}G_{\kappa}(LX,\mathrm{S}_{L}\phi+\psi)\cdot$ (206)
$\hskip 40.0pt\big{(}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}.$
In going to the last line, we used (108) to get $A^{-|Y|}\leqslant A^{-(1+\eta/2)|X|+9|\mathcal{C}(Y)|}$. Now for $\hat{K}_{j}$,
$\|\hat{K}_{j}^{(\cdot)}(\cdot,\psi)-\dot{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\|\hat{K}_{j}^{(\cdot)}(\cdot,\psi)-{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}+\|{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})-\dot{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}.$ (207)
The first term is more delicate. We use the fact that $\hat{K}_{j}$ is really only a function of $\mathrm{S}_{L}\phi+\psi$ to get
$\hat{K}_{j}^{X}$ $(\phi,\psi)-\hat{K}_{j}^{X}(\phi,\dot{\psi})$ (208)
$=\sum_{Y_{0}\in\mathcal{C}(X)}\int_{0}^{1}\prod_{Y\in\mathcal{C}(X)-{Y_{0}}}\tilde{K}_{j}(Y,\phi,(1-t)\psi+t\dot{\psi})\langle\mathrm{D}|_{\phi+t\mathrm{S}_{L^{-1}}(\dot{\psi}-\psi)}\tilde{K}_{j}(Y_{0},\cdot,\dot{\psi}),\mathrm{S}_{L^{-1}}(\dot{\psi}-\psi)\rangle\,\mathrm{d}t,$
so
$\|\hat{K}_{j}^{(\cdot)}(\cdot,\psi)-\hat{K}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}$ (209)
[…]
$\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-(1+\eta/3)|X|}\|\psi-\dot{\psi}\|_{\mathcal{C}^{2}(LX)}G_{\kappa}(LX,\mathrm{S}_{L}\phi+(1-t)\psi+t\dot{\psi})\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}^{|\mathcal{C}(X)|-1}\|K_{j},\dot{K}_{j}\|_{2\mathfrak{h},\kappa,A}.$
For the second term,
$\|{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})$ $-\dot{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\sum_{\varnothing\neq\mathcal{A}\subset\mathcal{C}(X)}\prod_{Y\in\mathcal{A}}\|\tilde{K}_{j}(Y,\cdot,\dot{\psi})-\dot{\tilde{K}}_{j}(Y,\cdot,\dot{\psi})\|_{\mathfrak{h};\langle Y,\phi\rangle}\prod_{Y\in\mathcal{A}^{c}}\|\dot{\tilde{K}}_{j}(Y,\cdot,\dot{\psi})\|_{\mathfrak{h};\langle Y,\phi\rangle}$
$\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-(1+\eta/3)|X|}G_{\kappa}(LX,\mathrm{S}_{L}\phi+\dot{\psi})\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}^{|\mathcal{C}(X)|-1}\cdot$ (210)
$\hskip 30.0pt\big{(}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}.$
By combining all this (we regroup some terms to isolate the ones which vanish when $\varepsilon=\dot{\varepsilon}$)
$\|\hat{K}_{j}^{(\cdot)}(\cdot,\psi)$ $-\dot{\hat{K}}_{j}^{(\cdot)}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle X,\phi\rangle}$
$\leqslant\mathcal{O}_{A}^{|\mathcal{C}(X)|}A^{-(1+\eta/3)|X|}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}^{|\mathcal{C}(X)|-1}\sup_{t\in[0,1]}G_{\kappa}(LX,\mathrm{S}_{L}\phi+(1-t)\psi+t\dot{\psi})\cdot$
$\hskip 20.0pt\Big{(}\big{(}\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\|\zeta_{0;j}-\dot{\zeta}_{0,j}\|_{\wp}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}$
$\hskip 30.0pt+\|K_{j},\dot{K}_{j}\|_{2\mathfrak{h},\kappa,A}\big{(}\|\psi-\dot{\psi}\|_{\mathcal{C}^{2}(LX)}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0,j}\|_{\wp}\big{)}\Big{)}.$ (211)
From here, the estimate on $R^{(\cdot)}_{j}$ follows immediately. Finally, we consider $P^{(\cdot)}_{j}$. For the case $\psi\neq\dot{\psi}$, a simple computation shows
$P_{j}^{X}$ $(\phi,\psi)-P_{j}^{X}(\phi,\dot{\psi})$ (212)
$=\sum_{B\in\mathcal{B}_{j+1}(X)}\int_{0}^{1}P_{j}^{X-\mathrm{Int}(B)}(\phi,(1-t)\psi+t\dot{\psi})e^{V_{j}(LB,\mathrm{S}_{L}\phi+(1-t)\psi+t\dot{\psi})}\frac{\mathrm{d}}{\mathrm{d}t}V_{j}(LB,\mathrm{S}_{L}\phi+(1-t)\psi+t\dot{\psi})\mathrm{d}t.$
Taking norms,
$\|P_{j}^{X}$ $(\cdot,\psi)-P_{j}^{X}(\cdot,\dot{\psi})\|_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(X)}$
$\leqslant\mathcal{O}_{L}\|\psi-\dot{\psi}\|_{\mathcal{C}^{2}(LX)}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+L^{-j}|\varepsilon-\dot{\varepsilon}|\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{)}\cdot$ (213)
$\hskip 40.0pt\sum_{B\in\mathcal{B}_{j+1}(X)}\|P_{j}^{(\cdot)}(\cdot,(1-t)\psi+t\dot{\psi})\|_{\mathfrak{h};\langle X-\mathrm{Int}(B),\phi\rangle}.$ (214)
Then for the difference,
$\|P_{j}^{(\cdot)}(\cdot,\psi)-\dot{P}_{j}^{(\cdot)}(\cdot,\psi)\|_{\mathfrak{h};\langle X,\phi\rangle}$
$\leqslant\bigg{\|}\prod_{B\in\mathcal{B}_{j+1}(X)}\big{(}e^{V_{j}(LB,\mathrm{S}_{L}(\cdot)+\psi)}-e^{V_{j+1}(B,\cdot)+\delta\mathcal{E}_{j}(LB)}\big{)}$
$\hskip 120.0pt-\prod_{B\in\mathcal{B}_{j+1}(X)}\big{(}e^{\dot{V}_{j}(LB,\mathrm{S}_{L}(\cdot)+\psi)}-e^{\dot{V}_{j+1}(B,\cdot)+\delta\dot{\mathcal{E}}_{j}(LB)}\big{)}\bigg{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(X)}$
$\leqslant\sum_{\varnothing\neq\mathcal{A}\subset\mathcal{B}_{j+1}(X)}\prod_{B\in\mathcal{A}^{c}}\|e^{\dot{V}_{j}(LB,\mathrm{S}_{L}(\cdot)+\psi)}-e^{\dot{V}_{j+1}(B,\cdot)+\delta\dot{\mathcal{E}}_{j}(LB)}\|_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(B)}\cdot$
$\hskip 40.0pt\prod_{B\in\mathcal{A}}\|e^{V_{j}(LB,\mathrm{S}_{L}(\cdot)+\psi)}-e^{V_{j+1}(B,\cdot)+\delta\mathcal{E}_{j}(LB)}-e^{\dot{V}_{j}(LB,\mathrm{S}_{L}(\cdot)+\psi)}+e^{\dot{V}_{j+1}(B,\cdot)+\delta\dot{\mathcal{E}}_{j}(LB)}\|_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(B)}$
$\leqslant\mathcal{O}_{L}^{|X|}\big{(}\|\zeta_{0;j},\dot{\zeta}_{j}\|_{\wp}+\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}\cdot$ (215)
$\hskip 20.0pt\Big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}-\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}+L^{-j}|\varepsilon-\dot{\varepsilon}|\big{(}\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j},\dot{K}_{j}\|_{\mathfrak{h},\kappa,A}\big{)}\Big{)}.$
Getting the last inequality requires some computations which should be more or less standard by now. Finally, we remark that the factors $\mathcal{O}_{L}$ which are generated can be used to absorb $\varrho^{(2)}_{\varepsilon;j}$ in the special case $j=-\mathsf{N}-1,\varepsilon=\dot{\varepsilon}$.
D.2 Proof of Proposition 2.3
We decompose
$K_{j+1}-\dot{K}_{j+1}=(K_{j+1}[\varepsilon,\zeta]-K_{j+1}[\varepsilon,\dot{\zeta}])+(K_{j+1}[\varepsilon,\dot{\zeta}]-K_{j+1}[\dot{\varepsilon},\dot{\zeta}]).$ (216)
For the first piece, we break the RG into the linear and nonlinear parts: $K_{j+1}=\mathcal{L}K_{j}+\mathcal{R}K_{j}$ where
$\mathcal{L}K_{j}$ $=\mathcal{L}^{(1)}K_{j}+\mathcal{L}^{(2)}K_{j}+\mathcal{L}^{(3)}K_{j},$
$\mathcal{R}K_{j}$ $=\mathcal{R}^{(1)}K_{j}+\mathcal{R}^{(2)}K_{j}+\mathcal{R}^{(3)}K_{j}+\mathcal{R}^{(4)}K_{j}+\mathcal{R}^{(5)}K_{j}+\mathcal{R}^{(6)}K_{j};$ (217)
and where
$\mathcal{L}^{(1)}K_{j}$ $=\sum_{Y\in\mathcal{S}_{j}}^{\bar{Y}^{L}=X}\mathrm{Rem}_{0}\mathbb{E}\big{[}k_{j}^{(0)}(Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})\big{]},$
$\mathcal{L}^{(2)}K_{j}$ $=\sum_{q\neq 0}\sum_{Y\in\mathcal{S}_{j}}^{\bar{Y}^{L}=X}\mathbb{E}\big{[}k_{j}^{(q)}(Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})\big{]},$
$\mathcal{L}^{(3)}K_{j}$ $=\sum_{Y\in\mathcal{P}_{\mathrm{c};j}-\mathcal{S}_{j}}^{\bar{Y}^{L}=X}\mathbb{E}\big{[}K_{j}(Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})\big{]},$
$\mathcal{R}^{(1)}K_{j}$ $=\mathbb{I}_{|X|=1}\big{(}\mathbb{E}\big{[}e^{V_{j}(LX-Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})}\big{]}-e^{V_{j+1}(X,\phi)+\delta\mathcal{E}_{j}(LX)}+\delta\mathcal{E}_{j}(LX)\big{)},$
$\mathcal{R}^{(2)}K_{j}$ $=\sum_{Y\in\mathcal{P}_{\mathrm{c};j}}^{\bar{Y}^{L}=X}\mathbb{E}\big{[}(e^{V_{j}(LX-Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})}-1)K_{j}(Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})\big{]},$
$\mathcal{R}^{(3)}K_{j}$ $=\sum_{Y\in\mathcal{P}_{j}}^{\bar{Y}^{L}=X,|\mathcal{C}(Y)|\geqslant 2}\mathbb{E}\bigg{[}e^{V_{j}(LX-Y,\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})}\prod_{Y^{\prime}\in\mathcal{C}(Y)}K_{j}(Y^{\prime},\mathrm{S}_{L}\phi+\Psi_{\hskip-1.0ptj})\bigg{]},$
$\mathcal{R}^{(4)}K_{j}$ $=\mathbb{I}_{|X|\geqslant 2}\,\mathbb{E}\big{[}P_{j}^{X}(\phi,\Psi_{\hskip-1.0ptj})\big{]}+(e^{-\delta\mathcal{E}_{j}(LX)}-1)\mathbb{E}\big{[}P_{j}^{X}(\phi,\Psi_{\hskip-1.0ptj})\big{]},$
$\mathcal{R}^{(5)}K_{j}$ $=\sum_{(Z,Y_{1},\mathfrak{B}\rightarrow Y_{2})\rightarrow X}^{|\mathcal{C}(Y_{1}\cup Y_{2})|\geqslant 1,|Z|+|\mathcal{C}(Y_{1}\cup Y_{2})|\geqslant 2}\mathbb{E}\big{[}P_{j}^{Z}(\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}},$ (218)
$\mathcal{R}^{(6)}K_{j}$ $=\sum_{(Z,Y_{1},\mathfrak{B}\rightarrow Y_{2})\rightarrow X}^{|\mathcal{C}(Y_{1}\cup Y_{2})|\geqslant 1}(e^{V_{j+1}(X-Z\cup Y_{1}\cup Y_{2},\phi)-\delta\mathcal{E}_{j}(L(Z\cup Y_{1}\cup Y_{2}))}-1)\mathbb{E}\big{[}P_{j}^{Z}(\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}.$
We first control the linear parts. For $\mathcal{L}^{(1)}K_{j}$, compute
$\|\mathcal{L}^{(1)}K_{j}[\varepsilon,\zeta]$ $-\mathcal{L}^{(1)}K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\sum_{Y\in\mathcal{S}}^{\bar{Y}^{L}=X}\big{\|}\mathrm{Rem}_{0}\mathbb{E}\big{[}(k_{j}^{(0)}[\varepsilon,\zeta]-k_{j}^{(0)}[\varepsilon,\dot{\zeta}])(Y,\mathrm{S}_{L}(\cdot)+\Psi_{j})\big{]}\big{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(\bar{Y}^{L})}$
$\lesssim_{1}L^{-2}\log(L)G_{\kappa}(X,\phi)\sum_{Y\in\mathcal{S}}^{\bar{Y}^{L}=X}\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa;\langle Y\rangle}$
$\lesssim_{1}A^{-|X|}\log(L)G_{\kappa}(X,\phi)\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa;\langle LX\rangle}$ (219)
For $\mathcal{L}^{(2)}K_{j}$,
$\|\mathcal{L}^{(2)}K_{j}[\varepsilon,\zeta]$ $-\mathcal{L}^{(2)}K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\sum_{q\neq 0}\sum_{Y\in\mathcal{S}}^{\bar{Y}^{L}=X}\big{\|}\mathbb{E}\big{[}(k_{j}^{(q)}[\varepsilon,\zeta]-k_{j}^{(q)}[\varepsilon,\dot{\zeta}])(Y,\mathrm{S}_{L}(\cdot)+\Psi_{j})\big{]}\big{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(\bar{Y}^{L})}$
$\lesssim_{1}L^{2}{e^{-\frac{1}{2}\beta\Gamma_{\hskip-1.0ptj}(0)}}A^{-|X|}G_{\kappa}(X,\phi)\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa;\langle LX\rangle}\sum_{n=0}^{\infty}e^{-(\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}\Gamma_{\hskip-1.0ptj}(0)-\mathfrak{h})\sqrt{\hskip-0.5pt\beta\hskip 0.5pt}n}$
$\lesssim_{1}L^{2\sigma}A^{-|X|}G_{\kappa}(X,\phi)\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa;\langle LX\rangle}.$ (220)
Finally, for $\mathcal{L}^{(3)}K_{j}$,
$\|\mathcal{L}^{(3)}K_{j}[\varepsilon,\zeta]$ $-\mathcal{L}^{(2)}K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\sum_{Y\in\mathcal{P}_{\mathrm{c};j}-\mathcal{S}_{j}}^{\bar{Y}^{L}=X}\big{\|}\mathbb{E}\big{[}(K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}])(Y,\mathrm{S}_{L}(\cdot)+\Psi_{j})\big{]}\big{\|}_{\mathfrak{h};\langle\phi\rangle;\mathcal{C}^{2}(\bar{Y}^{L})}$
$\leqslant\mathcal{O}_{L}^{|X|}A^{-(1+\eta)|X|}G_{\kappa}(X,\phi)\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa,A}.$ (221)
For the nonlinear terms, there is a bit of computation. We only treat one of the more complicated terms, since the rest are very similar or even easier. Precisely, we focus on $\mathcal{R}^{(6)}$. First we expand ($W=Z\cup Y_{1}\cup Y_{2}$ for brevity)
$(e$ $^{V_{j}[\varepsilon,\zeta](X-W,\phi)-\delta\mathcal{E}_{j}[\varepsilon,\zeta](LW)}-1)\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})R_{j}^{Y_{1}}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]$
$\hskip 5.0pt-(e^{V_{j}[\varepsilon,\dot{\zeta}](X-W,\phi)-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)}-1)\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\dot{\zeta}](\phi,\Psi_{\hskip-1.0ptj})R_{j}^{Y_{1}}[\varepsilon,\dot{\zeta}](\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\dot{\zeta}]$ (222)
$=e^{V_{j}[\varepsilon,\zeta](X-W,\phi)}(e^{-\delta\mathcal{E}_{j}[\varepsilon,\zeta](LW)}-e^{-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)})\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]+$
$\hskip 15.0pt(e^{V_{j}[\varepsilon,\zeta]}-e^{V_{j}[\varepsilon,\dot{\zeta}]})(X-W,\phi)e^{-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)}\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]+$
$\hskip 15.0pt(e^{V_{j}[\varepsilon,\dot{\zeta}](X-W,\phi)-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)}-1)\mathbb{E}\big{[}(P_{j}^{Z}[\varepsilon,\zeta]-P_{j}^{Z}[\varepsilon,\dot{\zeta}])(\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}[\varepsilon,\zeta](\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]+$
$\hskip 15.0pt(e^{V_{j}[\varepsilon,\dot{\zeta}](X-W,\phi)-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)}-1)\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\dot{\zeta}](\phi,\Psi_{\hskip-1.0ptj})(R^{Y_{1}}_{j}[\varepsilon,\zeta]-R^{Y_{1}}_{j}[\varepsilon,\dot{\zeta}])(\phi,\Psi_{\hskip-1.0ptj})\big{]}J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]+$
$\hskip 15.0pt(e^{V_{j}[\varepsilon,\dot{\zeta}](X-W,\phi)-\delta\mathcal{E}_{j}[\varepsilon,\dot{\zeta}](LW)}-1)\mathbb{E}\big{[}P_{j}^{Z}[\varepsilon,\dot{\zeta}](\phi,\Psi_{\hskip-1.0ptj})R^{Y_{1}}_{j}[\varepsilon,\dot{\zeta}](\phi,\Psi_{\hskip-1.0ptj})\big{]}(J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\zeta]-J_{j}^{\mathfrak{B}\rightarrow Y_{2}}[\varepsilon,\dot{\zeta}]).$
Then taking norms,
$\|\!\cdot\!\|_{\mathfrak{h};\langle X,\phi\rangle}\leqslant\mathcal{O}_{L}^{|X|}$ $\mathcal{O}_{A}A^{-(1+\eta/4)|X|}G_{\kappa}(X,\phi)\cdot$
$\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}\big{(}\|\zeta_{0;j}-\dot{\zeta}_{0;j}\|_{\wp}+\|K_{j}[\varepsilon,\zeta]-K_{j}[\varepsilon,\dot{\zeta}]\|_{\mathfrak{h},\kappa,A}\big{)}.$ (223)
Summing gives the kind of term required for the first part.
For the second part, we proceed similarly, but now there is no need to isolate the linear terms. Hence we take the basic definition of $K_{j+1}$, write down an identity like (222), and find a bound by the quantity
$L^{-j}|\varepsilon-\dot{\varepsilon}|\mathcal{O}_{L}^{|X|}A^{-(1+\eta/3)|Y_{1}\cup Y_{2}^{*}|}G_{2\kappa}(Y_{1},\phi)\|\zeta_{0;j},\dot{\zeta}_{0;j}\|_{\wp}^{|Z|}\big{(}\mathcal{O}_{A}\|K_{j}[\varepsilon,\dot{\zeta}],K_{j}[\dot{\varepsilon},\dot{\zeta}]\|_{2\mathfrak{h},2\kappa,A}\big{)}^{\!|\mathcal{C}(Y_{1}\cup Y_{2})|}.$ (224)
Ultimately, by summing over sets, the proposition follows.
A final remark as regards $j=-\mathsf{N}-1$: since
$K_{-\mathsf{N}}(X,\phi)=\mathbb{E}\big{[}P^{X}_{-\mathsf{N}-1}(\phi,\Psi_{\hskip-1.0pt-\mathsf{N}-1})\big{]},$ (225)
much simpler computations than above verify this special case.
Appendix E Laplace transforms
We are considering probability measures $\mu$ on $\mathcal{D}^{\prime}(\mathbb{T}^{d})$ for which the Laplace transform
$\mathrm{Z}_{\mu}(f)=\int e^{\langle\phi,f\rangle}\mu(\mathrm{d}\phi)$ (226)
belongs to $\mathfrak{C}$. In this case the complex Laplace transform
$\mathrm{Z}_{\mu}^{\mathbb{C}}(f)=\int e^{\langle\phi,\mathrm{Re}(f)\rangle+i\langle\phi,\mathrm{Im}(f)\rangle}\mu(\mathrm{d}\phi)$ (227)
is still Fréchet-analytic on $\mathcal{C}^{\infty}(\mathbb{T}^{d};\mathbb{C})$. We want to understand how uniform convergence of Laplace transforms as continuous functions gives us information about weak convergence of measures. As far as general notions of infinite-dimensional complex analysis go, [Din99] is a comprehensive reference.
Proposition E.1.
Suppose $\mu_{n}$ is a sequence of probability measures on $\mathcal{D}^{\prime}(\mathbb{T}^{d})$ with $\mathrm{Z}_{\mu_{n}}\in\mathfrak{C}$. Suppose further that $\mathrm{Z}_{\mu_{n}}$ converge in norm to some functional $\mathrm{Z}$. Then $\mathrm{Z}$ is the Laplace transform of a probability measure $\mu$ with Laplace transform in $\mathfrak{C}$, $\mu_{n}\rightarrow\mu$ weakly, and $\mathrm{Z}_{\mu}=\mathrm{Z}$.
Proof.
Consider some vectors $f_{1},\ldots,f_{r},g_{1},\ldots,g_{r}\in\mathcal{C}^{\infty}$. Define
$F_{n}(z_{1},\ldots,z_{2r})=\mathrm{Z}^{\mathbb{C}}_{\mu_{n}}(z_{1}f_{1}+\ldots+z_{r}f_{r}+z_{r+1}g_{1}+\ldots+z_{2r}g_{r}).$ (228)
These are holomorphic functions on $\mathbb{C}^{2r}$. On the real slice $\mathbb{R}^{2r}$, we know that $F_{n}$ converge locally uniformly to some $G$ defined by $\mathrm{Z}$. Now we want to show that every subsequence of $F_{n}$ contains a subsubsequence which converges to some $F$ (independent of the subsequence) and this will establish convergence.
Any subsequence $F_{n_{k}}$ does contain a subsubsequence converging to some holomorphic $F$ on $\mathbb{C}^{2r}$ by compactness (since $F_{n}$ are locally bounded and hence a normal family). Suppose that some other subsubsequence converges to an $F^{\prime}$ which might be different from $F$. But $F$ and $F^{\prime}$ coincide on $\mathbb{R}^{2r}$ and hence by the identity principle $F=F^{\prime}$.
With this established, it is immediate that $\mathrm{Z}^{\mathbb{C}}_{\mu_{n}}$ converge pointwise to some function $\mathrm{Z}^{\mathbb{C}}$ bounded by $G_{2\kappa_{*}}$ which extends $\mathrm{Z}$. It is also clear that it is Gâteaux-holomorphic:
$\mathrm{Z}^{\mathbb{C}}(u+iv+z_{1}(f_{1}+ig_{1})+\ldots+(f_{r}+ig_{r}))=\lim_{n\rightarrow\infty}F_{n}(1,i,z_{1},\ldots,z_{r},z_{1},\ldots,z_{r}),$ (229)
which is a uniform-on-compacts limit of restrictions of holomorphic functions to affine subspaces. Therefore the Gâteaux derivatives indeed exist. But a locally bounded Gâteaux-holomorphic function is Fréchet-holomorphic. In particular, this means that the characteristic functions $\mathrm{Z}^{\mathbb{C}}_{\mu_{n}}(if)$ converge pointwise to a function continuous at $x=0$, so by the Lévy-Fernique theorem for nuclear spaces (cf. e.g. [BDW17], Theorem 5.1 and references), $\mu_{n}$ converge weakly to a measure $\mu$ with characteristic function $\mathrm{Z}^{\mathbb{C}}(if)$. Since this characteristic function has a holomorphic extension, Cauchy bounds lead to moment estimates which suffice to conclude that $\mu$ indeed has a Laplace transform. Again by an identity principle, $\mathrm{Z}_{\mu}=\mathrm{Z}$. For completeness, we provide the details of this last step in the next proposition. ∎
Lemma E.2.
Suppose the characteristic function $\Phi$ of a probability measure $\mu$ on $\mathcal{D}^{\prime}(\mathbb{T}^{d})$ has a Gâteaux-holomorphic extension to $\mathcal{C}^{\infty}(\mathbb{T}^{d};\mathbb{C})$. Then this is the (rotated) Laplace transform of $\mu$.
Proof.
The main point is to establish moment bounds. Let $\mathrm{ev}_{\!f}$ by the evaluation functional at $f\in\mathcal{C}^{\infty}$, which becomes a random variable under $\mu$. Its characteristic function $\varphi_{\hskip-1.0ptf}(t)=\mu(e^{it\,\mathrm{ev}_{\!f}})$ satisfies
$\varphi_{\hskip-1.0ptf}(t)=\Phi(tf).$ (230)
Then $\varphi_{\hskip-1.0ptf}$ admits a holomorphic extension to $\mathbb{C}$. A well-known result ([Kle20], Theorem 15.35) states that since $\varphi_{\hskip-1.0ptf}$ is smooth,
$\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{2n}\big{]}\leqslant 2n|\varphi_{\hskip-1.0ptf}^{(2n)}(0)|.$ (231)
For odd moments we use
$\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{2n+1}\big{]}\leqslant 1+\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{2n+2}\big{]}\leqslant 1+(2n+2)|\varphi_{\hskip-1.0ptf}^{(2n+2)}(0)|.$ (232)
Integrating over a circle of radius $r$ gives
$\varphi_{\hskip-1.0ptf}^{(n)}(0)=\frac{n!}{2\pi i}\oint_{|z|=r}\frac{\varphi_{\hskip-1.0ptf}(z)}{z^{n+1}}\mathrm{d}z,$ (233)
that is,
$|\varphi_{\hskip-1.0ptf}^{(2n)}(0)|\leqslant(2n)!M_{f,r}r^{-2n}$ (234)
where $M_{f,r}$ is an upper bound on $\varphi_{\hskip-1.0ptf}$ on $|z|\leqslant r$. It follows that
$\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{2n}\big{]}$ $\leqslant 2n(2n)!M_{f,r}r^{-2n},$
$\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{2n+1}\big{]}$ $\leqslant 1+(2n+2)(2n+2)!M_{f,r}r^{-2n-2}.$ (235)
But we can take $r=2$ and then the integral
$\int e^{|\langle\phi,f\rangle|}\mu(\mathrm{d}\phi)=\sum_{k=0}^{\infty}\frac{1}{k!}\mathbb{E}\big{[}|\mathrm{ev}_{\!f}|^{k}\big{]}\leqslant e+M_{f,2}\sum_{k=0}^{\infty}(k+1)^{2}2^{-k}$ (236)
converges. Thus $\mu$ has a Laplace transform which has to be the given extension of $\Phi$ by the identity principle (applied, as before, to “finite-dimensional slices” of $\Phi$). ∎
Finally, we state a result about the tails of the measure given a Gaussian bound on the Laplace transform. The result below can probably be formulated more abstractly, but we provide a direct proof avoiding heavy machinery.
Proposition E.3.
Suppose $\mu$ is a measure on $\mathcal{D}^{\prime}(\mathbb{T}^{d})$ with Laplace transform bounded as
$\mathrm{Z}_{\mu}(f)\leqslant c\,e^{\mathsf{p}(f)^{\mathfrak{q}}}$ (237)
where $\mathsf{p}$ is some continuous seminorm on $\mathcal{C}^{\infty}(\mathbb{T}^{d})$, $c>0$, and $\mathfrak{q}\in(1,\infty)$ with Hölder conjugate $\mathfrak{p}=\frac{\mathfrak{q}}{\mathfrak{q}-1}$. Then there are an $\mathrm{L}^{\mathfrak{p}}$-based Sobolev norm $\|\!\cdot\!\|_{\natural}$ and $\lambda>0$ small such that
$\mu\big{(}e^{\lambda\|\cdot\|^{\mathfrak{p}}_{\natural}}\big{)}<\infty.$ (238)
Proof.
Let $\mathfrak{F}$ be the Fourier transform, which induces an isomorphism of the dual pairs $(\mathcal{C}^{\infty}(\mathbb{T}^{d}),\mathcal{D}^{\prime}(\mathbb{T}^{d}))$ and $(\mathcal{S}_{\mathrm{i}}^{\vphantom{a}}(\mathbb{Z}^{d};\mathbb{C}),\mathcal{S}_{\mathrm{i}}^{\prime}(\mathbb{Z}^{d};\mathbb{C}))$. The second couple are just function spaces on $\mathbb{Z}^{d}$ with the natural polynomial growth conditions, where the subscript indicates that the functions in question satisfy $f(-x)=\overline{f(x)}$. The duality pairing is explicitly given by
$\langle f,g\rangle_{\mathcal{S}^{\vphantom{a}}_{\mathrm{i}}\hskip-0.6pt(\mathbb{Z}^{d}),\mathcal{S}^{\prime}_{\mathrm{i}}\hskip-0.6pt(\mathbb{Z}^{d})}=\sum_{x\in\mathbb{Z}^{d}}\overline{\hskip-1.0ptf(x)\hskip-1.0pt}\,g(x).$ (239)
A natural system of norms is given by ($p\in[1,\infty],s\in\mathbb{R}$)
$\|f\|_{s,p}=\|(1+|\cdot|_{\infty})^{s}f\|_{\ell^{p}(\mathbb{Z}^{d};\mathbb{C})}.$ (240)
If we consider the pushforward measure $\mu_{*}=\mathfrak{F}_{*}\mu$, we find
$\mathrm{Z}_{\mu_{*}}\!(f)=\int e^{\langle f,\mathfrak{F}\phi\rangle}\mu(\mathrm{d}\phi)=\int e^{\langle\mathfrak{F}^{-1}\!f,\phi\rangle}\mu(\mathrm{d}\phi)=\mathrm{Z}_{\mu}(\mathfrak{F}^{-1}\hskip-1.0ptf).$ (241)
Thus, there are $s\in\mathbb{R}$, $\alpha>0$ such that
$\mathrm{Z}_{\mu_{*}}\!(f)\leqslant c\,e^{\alpha\|f\|_{s,\mathfrak{q}}^{\mathfrak{q}}/\mathfrak{q}}.$ (242)
If we define
$\delta_{x}^{\mathrm{Re}}=\frac{\delta_{x}+\delta_{-x}}{2},\quad\delta_{x}^{\mathrm{Im}}=\frac{\delta_{x}-\delta_{-x}}{2i},$ (243)
then
$|f(x)|^{p}\leqslant 2^{p-1}\big{(}|\langle\delta_{x}^{\mathrm{Re}},f\rangle|^{p}+|\langle\delta_{x}^{\mathrm{Im}},f\rangle|^{p}\big{)}$ (244)
and we can use Hölder’s inequality to compute
$\mu_{*}$ $\hskip-1.0pt\big{(}e^{\lambda\|\cdot\|_{-r,\mathfrak{p}}^{\mathfrak{p}}}\hskip-1.0pt\big{)}$
[…]
Apply Markov’s inequality with $t\geqslant 0$ arbitrary:
[…]
$\leqslant\inf_{y\in[0,\infty)}e^{-yt}\big{(}\mathrm{Z}_{\mu_{*}}\!(t\delta^{u}_{x})+\mathrm{Z}_{\mu_{*}}\!(-t\delta^{u}_{x})\big{)}$
$\leqslant 2c\inf_{y\in[0,\infty)}e^{2^{1-\mathfrak{q}}\alpha y^{\mathfrak{q}}(1+|x|_{\infty})^{s\mathfrak{q}}/\mathfrak{q}-yt}$
$=2c\,e^{-2\mathfrak{p}^{-1}\alpha^{-\mathfrak{p}/{\mathfrak{q}}}(1+|x|_{\infty})^{-s\mathfrak{p}}t^{\mathfrak{p}}},$ (246)
and therefore if $n\geqslant 1$,
[…]
Using this, it follows that if $r>s+d/\mathfrak{p}$ and $\lambda$ small enough,
$\mu_{*}\hskip-1.0pt\big{(}e^{\lambda\|\cdot\|_{-r,\mathfrak{p}}^{\mathfrak{p}}}\hskip-1.0pt\big{)}\leqslant c+2c\sum_{n=1}^{\infty}\bigg{(}2^{d}d\,\mathfrak{p}\alpha^{\mathfrak{p}/{\mathfrak{q}}}\lambda\sum_{k=1}^{\infty}k^{-(r-s)\mathfrak{p}+d-1}\bigg{)}^{\!n}<\infty.$ (248)
$\|\!\cdot\!\|_{\natural}=\|\mathfrak{F}(\cdot)\|_{-r,\mathfrak{p}}$ is the Sobolev norm we were after. ∎
Acknowledgments
I thank Roland Bauerschmidt, Jon Dimock, Massimiliano Gubinelli, Peter Paulovics, and Sarah-Jean Meyer for useful input. Special thanks go to David Brydges for his help and many enlightening conversations without which this paper might have appeared. The work was supported by Oxford’s Clarendon Fund.

### v2

@@
 The interested reader can easily modify the results herein to cover this case, which can anyway be treated by other methods giving stronger results.
-Furthermore pick $\varepsilon>0$ and then let $\gamma[C_{\varepsilon}]$ be the law of the mollified unit mass GFF $\chi_{\varepsilon}*\Phi$ (where $\chi_{\varepsilon}$ is a standard mollifier).
-[^3] We put in a mass to keep things as simple as possible, primarily to avoid issues with the zero mode which arise in the massless model.
+Furthermore pick $\varepsilon>0$ and then let $\gamma[C_{\varepsilon}]$ be the law of the mollified unit mass [^3] We put in a mass to keep things as simple as possible, primarily to avoid issues with the zero mode which arise in the massless model.
 Setting it equal to $1$ is convenient because it equals the width of the torus.
 Of course, this is not a “generic” choice, but the additional complications in case the mass is not fixed to the torus size did not seem illuminating enough to warrant investigation in this note.
+GFF $\chi_{\varepsilon}*\Phi$ ($\chi_{\varepsilon}$ is a not-too-smooth radially symmetric compactly supported mollifier with integral $1$, see Appendix A.1).
 Put
@@
 […]
-$\mu_{\mathrm{sG}}$ does make sense because $\mathrm{C}_{\varepsilon}$ is a UV-mollified GFF covariance with cutoff $\varepsilon$ so that the Gaussian measure $\gamma[\mathrm{C}_{\varepsilon}]$ is $\mathcal{C}^{\infty}$-valued.
+$\mu_{\mathrm{sG}}$ does make sense because $\mathrm{C}_{\varepsilon}$ is a UV-mollified GFF covariance with cutoff $\varepsilon$ so that the Gaussian measure $\gamma[\mathrm{C}_{\varepsilon}]$ is at least $\mathcal{C}^{0}$-valued.
 If $\mathrm{T}$ is a symmetry of the torus,
@@
 First, $\sigma\coloneqq 1-\frac{\beta}{8\pi}\in(0,\frac{1}{2}]$ is the distance to criticality.
+$\mathfrak{d}$ is an integer of order (say) $10$ and determines how many derivatives $\chi_{\varepsilon}$ has; equivalently, its decay in momentum space.
 $\delta$ is a small positive number which could be fixed at, say, $\delta=\frac{1}{100}$ throughout, $\ell\gtrsim_{\sigma^{-1},\delta^{-1}}1$ is an integer, $L=\ell^{\mathfrak{m}}$ where $\mathfrak{m}=\ell^{2}$, $A\gtrsim_{L}1$, and $\zeta_{*}\lesssim_{A^{-1}}1$.
@@
 […]
-Now we define
+This is legitimate because $\mathrm{C}_{\varepsilon}$ has polynomial decay in momentum space, hence any smooth $J$ is in its Cameron-Martin space. Now we define
 $\zeta_{\varepsilon;j}=L^{2j}e^{\frac{1}{2}\beta\sum_{k=j}^{0}\Gamma_{\hskip-1.0pt\varepsilon;k}(0)}\mathrm{S}_{L^{-j}}\zeta,\quad\text{i.e.}\quad\zeta_{\varepsilon;j+1}=L^{2}e^{-\frac{1}{2}\beta\Gamma_{\hskip-1.0pt\varepsilon;j}(0)}\mathrm{S}_{L^{-1}}\zeta_{\varepsilon;j}.$ (12)
@@
 Appendix A Scale decomposition
-A.1 Continuous decomposition
+A.1 Mollifier
+We take $\chi=c_{\mathfrak{d}}\mathsf{B}^{*2}\mathrm{C}^{*\mathfrak{d}}$ where $\mathsf{B}$ is a standard bump of support radius $1/4$ on $\mathbb{R}^{2}$ and $c_{\mathfrak{d}}$ is chosen so that $\chi$ has integral $1$.
+In Fourier space,
+$\hat{\chi}(p)\propto_{1}\int_{\mathbb{R}^{2}}\frac{\hat{\mathsf{B}}(p-q)^{2}}{(1+|q|^{2})^{\mathfrak{d}}}\mathrm{d}q.$ (79)
+Because $\hat{\mathsf{B}}$ is Schwartz, it is easy to see that this still behaves asymptotically like $(1+|p|^{2})^{\mathfrak{d}}$.
+For $0<\varepsilon<1$, put $\chi_{\varepsilon}(x)=\varepsilon^{-2}\chi(\varepsilon^{-1}x)$ (this also makes sense as a function on $\mathbb{T}^{2}$).
+A.2 Continuous decomposition
 Following [Bau13], we may realize the massive GFF on $\mathbb{T}^{2}$ as an integral against an $\mathrm{L}^{2}$-cylindrical Brownian motion $B_{t}$:
@@
 Here ${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5pt1)=0$ when $|x|_{2}\geqslant t/2$, and its precise form is
-${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5ptm)=t\sum_{p\in 2\pi\mathbb{Z}}\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}$ (80)
+${\mathrm{F}}_{\hskip-1.0ptt}(x;\hskip-0.5ptm)=t\sum_{p\in 2\pi\mathbb{Z}^{2}}\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}$ (81)
 where $\varphi$ is a real even $1/2$-Gevrey function [^5] We don’t need this much regularity, but we take it since it comes for free.
@@
 $\int_{0}^{\infty}t\,\varphi(t)^{2}\mathrm{d}t=1.$ (82)
-Let $\chi$ be the standard smooth bump function on $\mathbb{R}^{2}$ with support radius $1$.
-For $\varepsilon>0$, put $\chi_{\varepsilon}(x)=\varepsilon^{-2}\chi(\varepsilon^{-1}x)$.
 After mollifying and periodizing, we can decompose
 $\Phi_{\hskip-0.5pt\varepsilon}$ $=\int t^{-1/2}{\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(\cdot\,;\hskip-0.5ptm)*\mathrm{d}{B}_{t},$
-${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ $=(\chi_{\varepsilon}*{\mathrm{F}}_{\hskip-1.0ptt})(x)=t\hskip 1.0pt\sum_{p\in 2\pi\mathbb{T}^{2}}\hat{\chi}(\varepsilon p)\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}.$ (82)
+${\mathrm{F}}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ $=(\chi_{\varepsilon}*{\mathrm{F}}_{\hskip-1.0ptt})(x)=t\hskip 1.0pt\sum_{p\in 2\pi\mathbb{Z}^{2}}\hat{\chi}(\varepsilon p)\varphi\big{(}t\sqrt{m^{2}+p^{2}}\big{)}e^{ipx}.$ (83)
 The covariance decomposes as
 $\mathrm{C}_{\varepsilon}(x)=(\chi_{\varepsilon}*\mathrm{C}*\chi_{\varepsilon})(x)=\int_{0}^{\infty}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)\frac{\mathrm{d}t}{t}.$ (84)
-The following estimate is immediate for any multiindex $\alpha$ (it holds for $\varepsilon=0$ as well):
-$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}t^{-|\alpha|}e^{-\sqrt{tm}}.$ (84)
+The following estimate is immediate for any multiindex $\alpha\leqslant 4$ and $\varepsilon\geqslant 0$:
+$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{1}t^{-|\alpha|}e^{-\sqrt{tm}}.$ (85)
 When $t<\varepsilon$, we expect better estimates since we have mollified. Indeed, in this case
-$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{|\alpha|}(\varepsilon^{-1}t)^{2}\varepsilon^{-|\alpha|}.$ (85)
+$|\partial^{\alpha}_{x}{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)|\lesssim_{1}(\varepsilon^{-1}t)^{2}\varepsilon^{-|\alpha|}.$ (86)
 Finally, we wish to know the value at $x=0$ more accurately since it enters directly into the RG flow.
@@
 and the finite range property, $\mathrm{F}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)=\bar{\mathrm{F}}^{*2}_{\hskip-1.0pt\varepsilon;t}(x;\hskip-0.5ptm)$ if $t<1/4$, so for these values the above result transfers immediately.
-A.2 Discrete decomposition & scaling
+A.3 Discrete decomposition & scaling
 Write $\varepsilon=\epsilon L^{-\mathsf{N}-1}/6$ for $\epsilon\in[1,L)$. Then put ($j=-\mathsf{N},\ldots,-1$)
@@
 where $\varrho^{(1)}_{\varepsilon;j}$ is bounded uniformly away from $0,\infty$ if $j\neq-\mathsf{N}-1$. For the $j=0$ case, similar reasoning gives
-$|\partial^{\alpha}\Gamma_{\hskip-1.0pt\varepsilon;0}(x)|\lesssim_{|\alpha|}1.$ (94)
+$|\partial^{\alpha}\Gamma_{\hskip-1.0pt\varepsilon;0}(x)|\lesssim_{1}1.$ (95)
 Finally, we will also need to compare the fields at the same scales but with different covariances. Take $\dot{\varepsilon}\leqslant\varepsilon$ and restrict $j\neq-\mathsf{N}$. We have
-$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j})|(x)$ $\lesssim_{|\alpha|}L^{-j}|\varepsilon-\dot{\varepsilon}|,$ (95)
+$|\partial^{\alpha}(\Gamma_{\hskip-1.0pt\varepsilon;j}-\Gamma_{\hskip-1.0pt\dot{\varepsilon};j})|(x)$ $\lesssim_{1}L^{-j}|\varepsilon-\dot{\varepsilon}|,$ (96)
 $\mathbb{E}\big{[}\|\Psi_{\hskip-1.0pt\varepsilon;j}-\Psi_{\hskip-1.0pt\dot{\varepsilon};j}\|^{2}_{\mathcal{C}^{r}(X)}\big{]}$ $\lesssim_{r}L^{-2j}|\varepsilon-\dot{\varepsilon}|^{2}|X|,$ (97)
@@
 Showing all these relations is straightforward, using only Gaussian random variable theory, manipulations with the mollifier $\chi_{\varepsilon}$ in Fourier space, and the already stated properties of $\Gamma_{\hskip-1.0pt\varepsilon;j}$.
-A.3 Finer scale decomposition
+A.4 Finer scale decomposition
 We will need a refined scale decomposition to prove some estimates below.
@@
 $\langle\mathrm{D}^{n}|_{x}F,u_{1},\ldots,u_{n}\rangle=\sum_{m=0}^{\infty}\frac{1}{m!}\langle\mathrm{D}^{n+m}|_{x^{\prime}}F,x-x^{\prime},\ldots,x-x^{\prime},u_{1},\ldots,u_{n}\rangle.$ (130)
-Hence
+Hence (the series converge absolutely)
 $^{\mathbb{C\hskip-1.0pt}}F(x+z)$ $=\sum_{n,m=0}^{\infty}\frac{1}{n!m!}\langle\mathrm{D}^{n+m}|_{x^{\prime}}F,x-x^{\prime},\ldots,x-x^{\prime},z,\ldots,z\rangle$
@@
 $\leqslant\log G_{1}(Y,\phi)$ (163)
-where we have had to require $\mathfrak{c}^{(1)}\gtrsim_{1}\mathfrak{c}^{(2)}$. Moving on, by (135,140) and the triangle inequality,
+where we have had to require $\mathfrak{c}^{(1)}\gtrsim_{1}\mathfrak{c}^{(2)}$. Moving on, by (136,141) and the triangle inequality,
 $\log G_{1}(X,\mathrm{S}_{\ell}\phi+\psi)\leqslant\,$ $\mathfrak{c}^{(1)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}\alpha\mathfrak{c}^{(1)}\big{(}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(\partial X)}+\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}\big{)}+$
@@
 […]
-Now if $X$ is small and the conditions on constants as above, take (163) and combine it with (139,145):
+Now if $X$ is small and the conditions on constants as above, take (164) and combine it with (140,146):
 $2\log G_{1}(X,\mathrm{S}_{\ell}\phi+\psi)\leqslant\,$ $2\mathfrak{c}^{(1)}\|\nabla\mathrm{S}_{\ell}\phi\|^{2}_{\mathrm{L}^{2}(X)}+\mathcal{O}_{1}(\alpha+1)\mathfrak{c}^{(1)}\|\nabla^{2}\mathrm{S}_{\ell}\phi\|^{2}_{\mathcal{B}^{2}_{\infty}(X)}+$
@@
 For (168), first take $|\mu|\leqslant 4$ and estimate
-$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)}\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})).$ (172)
+$\mathbb{E}\big{[}\!\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\|\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s}\|^{2}_{\mathrm{L}^{2}(X)})\big{]}\leqslant\exp(\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Tr}\,\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})).$ (173)
 This holds as long as the $\mathrm{L}^{2}$ operator norm of $\mathcal{O}_{1}\kappa_{*}^{-1}\mathfrak{c}^{(3)}\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})$ is bounded by $1/2$, that is, we have to take $\kappa_{*}$ such that $\mathfrak{c}^{(3)}\|\mathrm{Cov}(\partial^{\mu}\Psi_{\hskip-1.0pt\varepsilon;j;s})\|_{\mathcal{B}(\mathrm{L}^{2}(X))}\lesssim_{1}\kappa_{*}$.
@@
 By assumption, $\mathrm{Rem}_{0}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}=\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(X,\mathrm{S}_{L}(\cdot)+\Psi_{\hskip-1.0pt\varepsilon;j})\big{]}$.
-From (185,186),
+From (186,187),
 $\big{\|}\mathrm{Rem}_{1}\mathbb{E}\big{[}k^{(0)}_{j}(\cdot,\mathrm{S}_{L}(\cdot)+\psi)\big{]}\big{\|}_{\mathfrak{h};\langle X,\phi\rangle;\mathcal{C}^{2}(\bar{X}^{L})}$ $\leqslant\|\mathrm{Rem}_{1}k^{(0)}_{j}\|_{8\mathfrak{h}/L;\langle X,\mathrm{S}_{L}\delta\phi+\psi\rangle}$
@@
 Let $\mathfrak{F}$ be the Fourier transform, which induces an isomorphism of the dual pairs $(\mathcal{C}^{\infty}(\mathbb{T}^{d}),\mathcal{D}^{\prime}(\mathbb{T}^{d}))$ and $(\mathcal{S}_{\mathrm{i}}^{\vphantom{a}}(\mathbb{Z}^{d};\mathbb{C}),\mathcal{S}_{\mathrm{i}}^{\prime}(\mathbb{Z}^{d};\mathbb{C}))$.
-The second couple are just function spaces on $\mathbb{Z}^{d}$ with the natural polynomial growth conditions, where the subscript indicates that the functions in question satisfy $f(-x)=\overline{f(x)}$.
+The second couple are just function spaces on $\mathbb{Z}^{d}$ with the natural polynomial growth conditions, where the subscript indicates that the functions in question satisfy $f(-x)=\overline{\hskip-1.0ptf(x)\hskip-1.0pt}\,$.
 The duality pairing is explicitly given by
@@
 I thank Roland Bauerschmidt, Jon Dimock, Massimiliano Gubinelli, Peter Paulovics, and Sarah-Jean Meyer for useful input.
-Special thanks go to David Brydges for his help and many enlightening conversations without which this paper might have appeared.
+Special thanks go to David Brydges for his help and many enlightening conversations without which this paper might not have appeared.
 The work was supported by Oxford’s Clarendon Fund.